### v2

# Existence and Uniqueness Results for a Mean-field Game of Optimal Investment

Abstract.
We establish the existence and uniqueness of the equilibrium for a stochastic mean-field game of optimal investment. The analysis covers both finite and infinite time horizons, and the mean-field interaction of the representative company with a mass of identical and indistinguishable firms is modeled through the time-dependent price at which the produced good is sold. At equilibrium, this price is given in terms of a nonlinear function of the expected (optimally controlled) production capacity of the representative company at each time. The proof of the existence and uniqueness of the mean-field equilibrium relies on a priori estimates and the study of nonlinear integral equations, but employs different techniques for the finite and infinite horizon cases. Additionally, we investigate the deterministic counterpart of the mean-field game under study.
a University of Parma, Department of Economics and Management, Via J. F. Kennedy 6, 43125 Parma (Italy).
b University of Bologna, Department of Mathematics, Piazza di Porta San Donato 5, 40126 Bologna (Italy).
c Bielefeld University, Center for Mathematical Economics (IMW), Universitätsstrasse 25, 33615 Bielefeld (Germany).
d LUISS University, Department of Economics and Finance, Viale Romania 32, 00197 Roma (Italy).
1 E-mail: alessandro.calvia@unipr.it.
2 E-mail: salvatore.federico@unibo.it.
3 E-mail: giorgio.ferrari@uni-bielefeld.de.
4 E-mail: fgozzi@luiss.it.
Giorgio Ferrari gratefully acknowledges financial support from Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)– Project-ID 317210226– SFB 1283. This work started during the visit of Alessandro Calvia, Salvatore Federico and Fausto Gozzi at the Center for Mathematical Economics (IMW) at Bielefeld University. Those authors thank IMW and the SFB 1283 for the support and hospitality.
Alessandro Calvia, Salvatore Federico and Fausto Gozzi are supported by the Italian Ministry of University and Research (MUR), in the framework of PRIN project 2017FKHBA8 001 (The Time-Space Evolution of Economic Activities: Mathematical Models and Empirical Applications).
Alessandro Calvia and Salvatore Federico are members of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica "Francesco Severi" (INdAM).
Keywords: mean-field games; mean-field equilibrium; forward-backward ODEs; optimal investment; price formation.
AMS 2020: 35Q89; 47H10; 49N10; 49N80; 91A07; 91B38; 91B70.
JEL classification: C02; C61; C62; C72; D25; D41.
1. Introduction
In this paper, we provide existence, uniqueness, and characterization results for the equilibrium of a mean-field model of optimal investment with competition à la Cournot. In the absence of interventions, the production capacity of the representative company evolves stochastically over time as a geometric Brownian motion, and its level can be increased through investment, subject to quadratic costs. The representative company discounts profits and costs at a constant rate and aims to maximize total expected profits from production, net of investment costs. Instantaneous profits depend linearly on the company’s production capacity (thus, production occurs at full capacity) and on the time-dependent price of the produced good. The mean-field equilibrium investment and average production processes $(\widehat{\bf{u}},\widehat{q})$ are such that expected total net profits are maximized and, assuming an isoelastic inverse demand function, the price is given in terms of a nonlinear function of the average optimally controlled production at each time (see also Achdou et al. [2] and Section 2.1 below). We are able to prove the existence and uniqueness of the equilibrium pair $(\widehat{\bf{u}},\widehat{q})$ when the problem’s time horizon is finite or infinite. These equilibria are characterized as the unique solution to a nonstandard system of forward-backward ordinary differential equations. Furthermore, we demonstrate that the existence and uniqueness carry over to the deterministic counterpart of our model, essentially using the same techniques as in the stochastic case.
The investment problem under study falls under the category of mean-field games with scalar interaction. Mean-field games, independently introduced by Lasry and Lions [18], and Huang, Caines, and Malhamé [16], represent limit models of non-cooperative symmetric $N$-player games with mean-field interaction as the number of players $N$ tends to infinity. An exhaustive review of mean-field models can be found in the two-volume book by Carmona and Delarue [8].
In our context, the consistency condition that the equilibrium price aligns at each time with a decreasing nonlinear function of the expected (optimally controlled) production capacity can be viewed as the limit, as the number $N$ of identical and indistinguishable companies operating in the market diverges, of the requirement that price inversely depends on the aggregate production of the entire economy, scaled by a factor of $1/N$. As discussed in Huang et al. [16], this scaling can be justified by considering situations where “an increasing number of firms join together to serve an increasing number of consumers” (see the discussion in Huang et al. [16], after Equation (2.4) therein).
The equilibrium construction for a given time horizon $T\in(0,+\infty]$ follows a three-step approach. Firstly, given a deterministic path of the average production $q\coloneqq(q_{s})_{s\in[0,T]}$, we solve the representative company’s optimal investment problem. This is of linear-quadratic type and has an explicit solution, since the production capacity evolves as a geometric Brownian motion and is linearly dependent on the investment process, and because the performance criterion has linear dependence on the production capacity and quadratic dependence on the investment cost. Moreover, the optimal control $\widehat{\mathbf{u}}^{(T,q)}$ is deterministic, for any fixed average production trajectory $q$. In the second step, we calculate the expectation of the optimally controlled production capacity process, which is easily computable given the explicit representation of the state process. Finally, in the third step, we impose the consistency condition, requiring that this expectation must match $q_{s}$ for each time $s\in[0,T]$. This condition leads to a nonlinear integral equation for the equilibrium average production trajectory $\widehat{q}\coloneqq(\widehat{q}_{s})_{s\in[0,T]}$. By construction the control $\widehat{\mathbf{u}}\coloneqq\widehat{\mathbf{u}}^{(T,\widehat{q})}$ and the function $\widehat{q}$ form a mean-field equilibrium.
It is important to note that although the company’s optimal control problem for a fixed average production trajectory $q$ is linear-quadratic, the overall mean-field problem is not of linear-quadratic type, due to the nonlinear dependence of the net profit functional on $q$. In this regard, our results differ from those presented in Bensoussan et al. [3], Delarue and Tchuendom [12], Tchuendom [20], among others, which focus on linear-quadratic mean-field games.
While the three-step approach previously outlined works for the case where $T<+\infty$ as well as the case where $T=+\infty$, different technical arguments are used to prove the existence and uniqueness of the equilibrium in the finite and infinite horizon cases. As a matter of fact, the integral equation that uniquely characterizes the equilibrium average production $\widehat{q}$ does not fall into the standard theory of integro-differential equations, as we have an initial-value problem with the integral being backward in time, and this requires a careful ad hoc analysis. Specifically, the proof of the existence of a solution relies on a priori estimates and Schauder’s fixed point theorem when $T<+\infty$, whereas it utilizes a priori estimates and the Frechet-Kolmorogorov ${\mathrm{L}}^{p}$ compactness theorem in the case of $T=+\infty$. Uniqueness is then established in both cases through suitable contradiction arguments that exploit the properties of mean-field equilibria. Notably, when $T=+\infty$, we can also prove that the time-dependent equilibrium $\widehat{q}$ monotonically converges to the stationary average production level. The latter is obtained as the unique constant solution to the integral equation that uniquely determines $\widehat{q}$.
Given the geometric dynamics of the production capacity, the linear dependence of the profit functional on the controlled state variable, the quadratic costs of investment, and the fact that the mean-field interaction is of scalar type and only involves the expected value of the latter, it is not surprising that the mean-field equilibrium is deterministic and does not depend on the volatility coefficient $\sigma$ appearing in the production capacity’s dynamics. Guided by this observation, we also consider the deterministic counterpart of the previously discussed mean-field game and show that a unique equilibrium exists in this setting as well. This equilibrium can indeed be constructed by following exactly the same arguments employed in the stochastic case.
Games of optimal investment in both stochastic and deterministic settings are extensively covered in the Economics literature, and a comprehensive overview of models and results can be found in Vives [21]. Specifically, mean-field problems with Cournot competition have garnered interest in recent literature; see Chan and Sircar [10] for an insightful overview. Graber and Bensoussan [15] study the existence (and, under certain conditions, uniqueness) of a solution to a system of partial differential equations (PDEs) (namely, the Hamilton-Jacobi-Bellman (HJB) and Fokker-Planck equations) associated with a mean-field game involving Bertrand and Cournot competition among a continuum of players. In Graber and Sircar [14], existence and uniqueness of the master equation associated with a mean-field game of controls with absorption are proven, while Chan and Sircar [9] delve into dynamic mean-field games with exhaustible capacities and interactions akin to Cournot and Bertrand competitions. An optimal transport perspective on Cournot-Nash equilibria is explored in Acciaio et al. [1]. Finally, Cao et al. [6] considers stationary discounted and ergodic mean-field games of singular controls motivated by irreversible investment and provide existence and uniqueness results, as well as relations across the two classes of considered problems.
To the best of our knowledge, the existence and uniqueness of the (nonstationary) mean-field equilibrium for a mean-field model of optimal investment with an isoelastic demand function, as discussed in this paper, is presented here for the first time. It is also worth noting that some of our results – in particular the existence and uniqueness ones – carry over also to the case of a more general instantaneous profit function and more general mean-field interaction, upon requiring suitable monotonicity conditions on the problem’s data. We refer to Section 2.1 below for further details.
The rest of the paper is organized as follows. The stochastic mean-field game is introduced in Section 2, and in Subsection 2.1 we provide its economic foundation. In Section 3 the company’s optimal investment problem is solved for both the cases $T<+\infty$ and $T=+\infty$, while existence and uniqueness of the equilibrium is shown in Section 4, again for the finite and infinite time horizon cases. The deterministic version of the mean-field problem is finally considered in Section 5.
1.1. Notation
In this section we collect the main notation used in this work.
  • Throughout the paper the set $\mathbb{N}$ denotes the set of natural integers without the zero element, i.e., =$\mathbb{N}=\{1,2,\dots\}$, while $\mathbb{R}$ denotes the set of real numbers. Whenever $T=+\infty$, the notation $[0,T]$ indicates the interval $[0,+\infty).$
  • For any $p\geq 1$, any measure space $(E,\mathscr{E},\mu)$, and any interval $I\subseteq\mathbb{R}$, we indicate by ${\mathrm{L}}^{p}((E,\mathscr{E},\mu);I)$ the set of all functions with values in $I$ that are $p$-integrable with respect to $\mu$. If $E\subseteq\mathbb{R}$, we take $\mathscr{L}(E)$ as the Lebesgue $\sigma$-algebra and as $\mu$ the Lebesgue measure, which is denoted by $\mathrm{Leb}$, and we simply write ${\mathrm{L}}^{p}(E;I)$. When computing integrals with respect to this measure, we simply write ${\mathrm{d}}x$ instead of $\mathrm{Leb}({\mathrm{d}}x)$.
  • For any $\eta>0$ and $E,\,I\subseteq\mathbb{R}$, the notation ${\mathrm{L}}^{p}_{\eta}(E;I)$ indicates the set of all functions $f\colon E\to I$ such that $t\mapsto{\mathrm{e}}^{-\eta t}f(t)$ is $p$-integrable with respect to the Lebesgue measure.
  • Given two intervals $I,J\subseteq\mathbb{R}$, we denote by ${\mathrm{C}}(I;J)$ the set of all continuous functions from $I$ to $J$, endowed with the usual sup-norm. The notation ${\mathrm{C}}^{p}(I;J)$, $p\in\mathbb{N}\cup\{+\infty\}$, denotes the set of functions from $I$ to $J$ that are continuously differentiable $p$ times.
2. The stochastic model
Let $(\Omega,{\mathcal{F}},{\mathbb{F}}\coloneqq({\mathcal{F}}_{s})_{s\geq 0},{\mathbb{P}})$ be a complete filtered probability space, with ${\mathbb{F}}$ satisfying the usual assumptions, supporting an ${\mathbb{F}}$-adapted standard Brownian motion $B$ and an independent ${\mathcal{F}}_{0}$-measurable random variable $\xi$. Throughout the paper, $T\in(0,+\infty)\cup\{+\infty\}$ denotes a finite or infinite time horizon. In what follows, whenever $T=+\infty$ the notation $[0,T]$ indicates the interval $[0,+\infty)$.
We consider a real-valued ${\mathbb{F}}$-adapted process $X=(X_{s})_{s\in[0,T]}$, satisfying the stochastic differential equation (SDE)
[…]
where $\delta,\sigma>0$ are given coefficients, and the control process ${\mathbf{u}}\coloneqq(u_{s})_{s\in[0,T]}$ is chosen in either of the following two classes of admissible controls: if $T<+\infty$,
[…]
$\qquad{\mathbb{E}}\biggl{[}\int_{0}^{T}u_{s}^{2}\,{\mathrm{d}}s\biggr{]}<+\infty\biggr{\}}\,;$ (2.2)
if, instead, $T=+\infty$,
[…]
$\qquad{\mathbb{E}}\left[\int_{0}^{t}u_{s}\,{\mathrm{d}}s\right]<+\infty,\,{\mathbb{P}}\text{-a.s., }\forall t\geq 0,\text{ and }{\mathbb{E}}\biggl{[}\int_{0}^{+\infty}{\mathrm{e}}^{-\rho s}u_{s}^{2}\,{\mathrm{d}}s\biggr{]}<+\infty\biggr{\}}\,.$ (2.3)
Whenever necessary, to stress the dependence of the solution to (2.1) on the initial condition $\xi$ and on the control ${\mathbf{u}}$, we denote it by $X^{\xi,{\mathbf{u}}}$.
We observe that the solution to SDE (2.1) can be written explicitly (cf. [17, Problem 5.6.15]), for each ${\mathbf{u}}\in{\mathcal{U}}_{T}$, as
$X_{s}^{\xi,{\mathbf{u}}}=Y_{s}\xi+\int_{0}^{s}\dfrac{Y_{s}}{Y_{r}}u_{r}\,{\mathrm{d}}r,\quad s\in[0,T],$ (2.4)
where
$Y_{s}\coloneqq{\mathrm{e}}^{\sigma B_{s}-\left(\delta+\frac{\sigma^{2}}{2}\right)s},\quad s\in[0,T].$ (2.5)
In the mean-field game studied in this paper, the process $X$ describes the private state of a representative player. In particular, it models the evolution of her/his production capacity, which depreciates at a rate $\delta$ and can be increased by choosing the investment rate ${\mathbf{u}}\in{\mathcal{U}}_{T}$. Note that the statistical distribution of the state of the other players in the economy does not influence the production capacity level of the representative player.
Throughout the paper, we work under the following assumption.
Assumption 2.1.
The initial condition $\xi$ of SDE (2.1) is an ${\mathcal{F}}_{0}$-measurable, positive, and integrable random variable. More precisely, $\xi>0$, ${\mathbb{P}}$-a.s., and $0<{\mathbb{E}}[\xi]<+\infty$.
The assumption above ensures that, for any ${\mathbf{u}}\in{\mathcal{U}}_{T}$, $X_{s}^{\xi,{\mathbf{u}}}\geq 0$, ${\mathbb{P}}$-a.s., for all $s\in[0,T]$. Thus, the production capacity level of the representative agent is never negative. Moreover, it grants us the following result, which will be used later on.
Lemma 2.2.
Under Assumption 2.1 and for any ${\mathbf{u}}\in{\mathcal{U}}_{T}$, the unique solution to SDE (2.1) has finite first moment, given by
${\mathbb{E}}[X^{\xi,{\mathbf{u}}}_{s}]={\mathbb{E}}[\xi]{\mathrm{e}}^{-\delta s}+{\mathbb{E}}\left[\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}u_{r}\,{\mathrm{d}}r\right],\quad s\in[0,T].$ (2.6)
Proof.
Fix $\xi$ and ${\mathbf{u}}$ as above. Using the expression of $X^{\xi,{\mathbf{u}}}$ given in (2.4) we have that
${\mathbb{E}}[X^{\xi,{\mathbf{u}}}_{s}]={\mathbb{E}}[Y_{s}\xi]+{\mathbb{E}}\left[\int_{0}^{s}\dfrac{Y_{s}}{Y_{r}}u_{r}\right]\,{\mathrm{d}}r,\quad s\in[0,T].$
Since the random variable $Y_{s}\xi$ is non-negative and $Y_{s}$ is independent of ${\mathcal{F}}_{0}$, for all $s\in[0,T]$, we can directly compute the first summand
${\mathbb{E}}[Y_{s}\xi]={\mathbb{E}}[\xi]{\mathbb{E}}[Y_{s}]={\mathbb{E}}[\xi]{\mathrm{e}}^{-\delta s}.$
For the second summand, observe that the integrand is non-negative and that, for each fixed $s\in[0,T]$, the random variable $\frac{Y_{s}}{Y_{r}}$ is independent of ${\mathcal{F}}_{r}$, for all $r\in[0,s]$. Therefore, applying the Fubini-Tonelli theorem,
${\mathbb{E}}\left[\int_{0}^{s}\dfrac{Y_{s}}{Y_{r}}u_{r}\,{\mathrm{d}}r\right]=\int_{0}^{s}{\mathbb{E}}\left[\dfrac{Y_{s}}{Y_{r}}u_{r}\right]{\mathrm{d}}r=\int_{0}^{s}{\mathbb{E}}\left[\dfrac{Y_{s}}{Y_{r}}\right]{\mathbb{E}}[u_{r}]\,{\mathrm{d}}r={\mathbb{E}}\left[\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}u_{r}\,{\mathrm{d}}r\right],$
which is finite thanks to the assumptions on ${\mathbf{u}}$. Therefore, we get (2.6). ∎
In our mean-field game, the representative company’s instantaneous profit function is derived from the isoelastic demand function of Spence-Dixit-Stiglitz preferences (see Achdou et al. [2, p. 7, Footnote 5] and Luttmer [19]), as explained in Section 2.1 below. Moreover, the company faces a quadratic cost of investment. This leads to the expected net profit functional
$J_{T,q}(\xi,{\mathbf{u}})\coloneqq{\mathbb{E}}\left[\int_{0}^{T}{\mathrm{e}}^{-\rho s}\left\{X_{s}^{\xi,{\mathbf{u}}}q_{s}^{-\beta}-\dfrac{1}{2}u_{s}^{2}\right\}{\mathrm{d}}s\right]\,.$ (2.7)
Here, $\rho>0$ is a discount factor, $\beta>0$ is a fixed parameter, and $q=(q_{s})_{s\in[0,T]}$ is a given deterministic measurable function. At equilibrium, the function $q$ will identify with the average production capacity of the whole population of agents. This is formalized in the next definition.
Definition 2.3.
Fix a random variable $\xi$ under Assumption 2.1.
A pair $(\widehat{\mathbf{u}},\widehat{q})$, where $\widehat{\mathbf{u}}\in{\mathcal{U}}_{T}$ and $\widehat{q}\colon[0,T]\to(0,+\infty)$ is a measurable function, is an equilibrium of the mean-field game if
  1. (i)
    $J_{T,\widehat{q}}(\xi,\widehat{\mathbf{u}})\geq J_{T,\widehat{q}}(\xi,{\mathbf{u}})$, for all ${\mathbf{u}}\in{\mathcal{U}}_{T}$;
  2. (ii)
    $\widehat{q}_{s}={\mathbb{E}}[X_{s}^{\xi,\widehat{\mathbf{u}}}]$, for all $s\in[0,T]$.
In the next sections we show that there exists a unique equilibrium by adopting a classic fixed point approach. The first step is to solve an optimization problem for a given measurable function $q$, representing the evolution of the average production capacity level of the agents in the economy. We show that there exists a unique explicit and deterministic optimal control $\widehat{\mathbf{u}}\in{\mathcal{U}}_{T}$, depending on $q$, which provides us with the best response of the representative agent to the distribution of the states of the other agents in the economy, summarized by the average $q$.
The second step is to determine the optimally controlled dynamics of the production capacity level of the representative agent. Thanks to (2.6), we can compute explicitly ${\mathbb{E}}[X^{\xi,\widehat{\mathbf{u}}}_{s}]$, $s\in[0,T]$. Since the optimal control $\widehat{\mathbf{u}}$ depends on $q$, the fixed point argument follows from condition (ii) in Definition 2.3. We show that proving that there exists a unique equilibrium reduces to finding the unique solution of an integro-differential equation.
2.1. The economic foundation
To provide an economic foundation of our model, let us introduce the $N$-player game corresponding to the mean-field game introduced above.
We consider a finite number of firms $N$, whose production capacities evolve stochastically over time according to the following dynamics
[…]
where $B^{1},\dots,B^{n}$ are independent Wiener processes and $\xi^{1},\dots,\xi^{N}$ are i.i.d., ${\mathcal{F}}_{0}$-measurable, positive, and integrable random variables, independent from the Brownian motions $B^{i}$, $i=1,\dots,n$.
Each firm may increase its own production capacity, by investing at rate ${\mathbf{u}}^{i}=(u^{i}_{s})_{s\in[0,T]}\in{\mathcal{U}}_{T}$, and aims at maximizing the expected net profit functional
${\mathbb{E}}\Biggl{[}\int_{0}^{T}{\mathrm{e}}^{-\rho s}\Biggl{\{}X^{i}_{s}\biggl{(}\frac{1}{N}\sum_{j=1}^{N}X^{j}_{s}\biggr{)}^{-\beta}-\dfrac{1}{2}(u^{i}_{s})^{2}\Biggr{\}}{\mathrm{d}}s\Biggr{]}\,.$ (2.9)
Taking the limit as $N\to\infty$ and considering Nash equilibria, we expect that the dynamics (2.8) of each firm converges to the dynamics (2.1) of the representative player of the mean-field game, and that at equilibrium the functional (2.9) converges to $J_{T,\widehat{q}}$ given in (2.7), with $\widehat{q}$ satisfying the consistency condition of Definition 2.3-(ii).
We leave for future research the study of the $N$-player game and its convergence to the corresponding mean-field game, since this requires a detailed and careful analysis which is outside the scope of this paper, as we intend to focus on directly studying the mean-field game.
Now, instead, we want to focus on providing the precise derivation of the instantaneous profit function appearing in (2.9), i.e., the function
$(x_{1},\dots,x_{n})\mapsto x_{i}\biggl{(}\frac{1}{N}\sum_{j=1}^{N}x^{j}\biggr{)}^{-\beta},$ (2.10)
for any fixed $i=1,\dots,N$. This in turn justifies the form of the profit function $(x,p)\mapsto xp^{-\beta}$ appearing in (2.7). We follow an approach close to the one discussed in Achdou et al. [2, p. 7, Footnote 5].
Assume that each player $i=1,\dots,N$ in our economy is a firm producing a single good and facing the demand function
$\left(\frac{\pi_{i}}{P}\right)^{-\gamma},\quad\text{with }\gamma>1,$
where $\gamma$ is the elasticity of substitution, $\pi$ is the price set by firm $i$ and $P=\left(\frac{1}{N}\sum_{j=1}^{N}\pi_{j}^{1-\gamma}\right)^{1/(1-\gamma)}$ is a price index. This choice of the demand function is standard in economics and is justified by the Spence-Dixit-Stiglitz preferences (see, e.g., Dixit and Stiglitz [13]).
Each firm aims at maximizing the function
$\pi_{i}\mapsto\left\{\pi_{i}\left(\frac{\pi_{i}}{P}\right)^{-\gamma}-\frac{1}{x_{i}^{\delta}}\left(\frac{\pi_{i}}{P}\right)^{-\gamma}\right\},$
where $x_{i}>0$ is the productivity (or size) of firm $i$ and, hence, costs decrease with respect to productivity, according to the factor $x_{i}^{-\delta}$, where $\delta>0$ is a given parameter. We can consider each firm production capacity as a proxy for $x_{i}$ and, thus, this setting is coherent with the model introduced above.
It is easy to see that the maximum profit is
$\frac{1}{\gamma-1}\left(\frac{\gamma}{\gamma-1}\right)^{-\gamma}x_{i}^{\delta(\gamma-1)}P^{\gamma}\,,$ (2.11)
which is attained setting the price $\pi_{i}=\frac{\gamma}{(\gamma-1)}x_{i}^{-\delta}$.
This entails that $P=\frac{\gamma}{(\gamma-1)}\left(\frac{1}{N}\sum_{j=1}^{N}x_{j}^{\delta(\gamma-1)}\right)^{-1/(\gamma-1)}$, and hence, substituting this expression into (2.11), we get that the profit is
$\frac{1}{\gamma-1}x_{i}^{\delta(\gamma-1)}\left(\frac{1}{N}\sum_{j=1}^{N}x_{j}^{\delta(\gamma-1)}\right)^{-\frac{\gamma}{\gamma-1}}\,.$ (2.12)
Therefore, setting $\beta=\frac{\gamma}{\gamma-1}$, $\delta=\frac{1}{\gamma-1}$, and $p=\frac{1}{N}\sum_{j=1}^{N}x_{j}$, we get the profit function (2.10) appearing in our model, except for the constant $\frac{1}{\gamma-1}$ which does not affect the optimization problem.
The discussion above highlights that it is possible to consider more general forms of interaction between players in the profit function, which can be taken of the form
$\Phi(x)\,F\left(\frac{1}{N}\sum_{j=1}^{N}f(x_{j})\right),$ (2.13)
for some $\Phi,f\colon\mathbb{R}\to[0,+\infty)$ strictly increasing and $F\colon[0,+\infty)\to[0,+\infty)$, see, e.g., [6]. However, with the aim of obtaining an explicit characterization of the equilibrium, in this paper we focus on the particular case in which $\Phi$ and $f$ are the identity and $F(x)=x^{-\beta}$, as in (2.12). Nonetheless, it is worth noting that part of our subsequent results – in particular the existence ones – could be extended to the more general case of profit function (2.13), upon exploiting the monotonicity of functions $\Phi,F,$ and $f$.
3. The optimization problem
In this section we consider the optimization problem associated to the mean-field game described in Section 2. We show that, for each possible choice (in a suitable class of measurable functions) of the function $q$ appearing in the discounted net profit functional (2.7), the corresponding maximization problem has a unique solution and we compute explicitly the associated optimal control.
We divide our analysis into two subsections, the first devoted to the finite time horizon case, the second one to the infinite time horizon case.
3.1. The finite time horizon case
Let us consider the finite time horizon case, i.e., fix $T<+\infty$. For each fixed $q\colon[0,T]\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$, we consider the problem of maximizing the functional
$J_{T,q}(\xi,{\mathbf{u}})\coloneqq{\mathbb{E}}\left[\int_{0}^{T}{\mathrm{e}}^{-\rho s}\left\{X^{\xi,{\mathbf{u}}}_{s}q_{s}^{-\beta}-\dfrac{1}{2}u_{s}^{2}\right\}{\mathrm{d}}s\right],$ (3.1)
where $\xi$ satisfies Assumption 2.1, ${\mathbf{u}}$ is chosen in the class of admissible controls introduced in (2), and $X^{\xi,{\mathbf{u}}}$ is the unique solution of SDE (2.1). We also introduce the value function corresponding to the optimization problem, namely,
$V_{T,q}(\xi)\coloneqq\sup_{{\mathbf{u}}\in{\mathcal{U}}_{T}}J_{T,q}(\xi,{\mathbf{u}}),\quad\xi\in{\mathrm{L}}^{1}((\Omega,{\mathcal{F}}_{0},{\mathbb{P}});(0,+\infty)).$ (3.2)
In the next proposition we are going to show important properties of the functional $J_{T,q}$.
Proposition 3.1.
Fix a random variable $\xi$ satisfying Assumption 2.1, ${\mathbf{u}}\in{\mathcal{U}}_{T}$, and $q\colon[0,T]\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$. Then, the functional $J_{T,q}$ defined in (3.1) is finite and verifies
$J_{T,q}(\xi,{\mathbf{u}})={\mathbb{E}}[\xi]z^{(T,q)}_{0}+{\mathbb{E}}\left[\int_{0}^{T}{\mathrm{e}}^{-\rho s}\left\{z^{(T,q)}_{s}u_{s}-\dfrac{1}{2}u_{s}^{2}\right\}{\mathrm{d}}s\right],$ (3.3)
where $z^{(T,q)}\colon[0,T]\to[0,+\infty)$ is the deterministic function given by
$z^{(T,q)}_{s}\coloneqq\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta)(r-s)}q_{r}^{-\beta}\,{\mathrm{d}}r,\quad s\in[0,T].$ (3.4)
Moreover, if $\xi,\xi^{\prime}$ both verify Assumption 2.1 and are such that ${\mathbb{E}}[\xi]={\mathbb{E}}[\xi^{\prime}]$, then $J_{T,q}(\xi;{\mathbf{u}})=J_{T,q}(\xi^{\prime};{\mathbf{u}})$, for any ${\mathbf{u}}\in{\mathcal{U}}_{T}$.
Proof.
Fix $\xi\in{\mathrm{L}}^{1}((\Omega,{\mathcal{F}}_{0},{\mathbb{P}});(0,+\infty))$, and ${\mathbf{u}}\in{\mathcal{U}}_{T}$. Since the assumptions of Lemma 2.2 are verified, we can use (2.6) and apply the Fubini-Tonelli theorem (note that the integrand is non-negative) to get
${\mathbb{E}}\left[\int_{0}^{T}{\mathrm{e}}^{-\rho s}X^{\xi,{\mathbf{u}}}_{s}q_{s}^{-\beta}\,{\mathrm{d}}s\right]={\mathbb{E}}[\xi]\int_{0}^{T}{\mathrm{e}}^{-\rho s}{\mathrm{e}}^{-\delta s}q_{s}^{-\beta}\,{\mathrm{d}}s+{\mathbb{E}}\left[\int_{0}^{T}\int_{0}^{s}{\mathrm{e}}^{-\rho s}{\mathrm{e}}^{-\delta(s-r)}q_{s}^{-\beta}u_{r}\,{\mathrm{d}}r\,{\mathrm{d}}s\right].$
Clearly, the first summand of the last equality is finite, thanks to the assumptions on $\xi$ and $q$. Moreover,
${\mathbb{E}}[\xi]\int_{0}^{T}{\mathrm{e}}^{-\rho s}{\mathrm{e}}^{-\delta s}q_{s}^{-\beta}\,{\mathrm{d}}s={\mathbb{E}}[\xi]z^{(T,q)}_{0}.$
Also the second summand is finite, thanks to the assumptions on $q$ and ${\mathbf{u}}$. Exchanging the order of the two time integrals we get
${\mathbb{E}}\left[\int_{0}^{T}\int_{0}^{s}{\mathrm{e}}^{-\rho s}{\mathrm{e}}^{-\delta(s-r)}q_{s}^{-\beta}u_{r}\,{\mathrm{d}}r\,{\mathrm{d}}s\right]={\mathbb{E}}\left[\int_{0}^{T}{\mathrm{e}}^{\delta r}u_{r}\int_{r}^{T}{\mathrm{e}}^{-(\rho+\delta)s}q_{s}^{-\beta}\,{\mathrm{d}}s\,{\mathrm{d}}r\right]$
$=$ ${\mathbb{E}}\left[\int_{0}^{T}{\mathrm{e}}^{-\rho r}u_{r}\int_{r}^{T}{\mathrm{e}}^{-(\rho+\delta)(s-r)}q_{s}^{-\beta}\,{\mathrm{d}}s\,{\mathrm{d}}r\right]={\mathbb{E}}\left[\int_{0}^{T}{\mathrm{e}}^{-\rho r}z^{(T,q)}_{r}u_{r}\,{\mathrm{d}}r\right].$
It easily follows that $J_{T,q}$ is finite and that Equation (3.3) holds, which also entails the last statement of the Proposition. ∎
Remark 3.2.
The last statement in Proposition 3.1 implies that the value function $V_{T,q}$ depends on the initial condition $\xi$ only through its average. More precisely, if $\xi,\xi^{\prime}$ both verify Assumption 2.1 and are such that ${\mathbb{E}}[\xi]={\mathbb{E}}[\xi^{\prime}]$, then $V_{T,q}(\xi)=V_{T,q}(\xi^{\prime})$.
Thanks to the Proposition above we can find the optimal control for our optimization problem and this, as a byproduct, allows us to explicitly compute the value function.
Theorem 3.3.
Fix a random variable $\xi$ satisfying Assumption 2.1 and $q\colon[0,T]\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$. Then, $\widehat{\mathbf{u}}^{(T,q)}\coloneqq z^{(T,q)}\in{\mathcal{U}}_{T}$, where $z^{(T,q)}$ is the function defined in (3.4), is an optimal control for problem (3.2), which is deterministic and independent of $\xi$.
Moreover, $\widehat{\mathbf{u}}^{(T,q)}$ is essentially unique, i.e., if $\overline{\mathbf{u}}^{(T,q)}\in{\mathcal{U}}_{T}$ is an optimal control for problem (3.2) different from $\widehat{\mathbf{u}}^{(T,q)}$, then
$\overline{u}^{(T,q)}_{s}=\widehat{u}^{(T,q)}_{s},\quad\text{for ${\mathbb{P}}\otimes\mathrm{Leb}$-a.e. }(\omega,s)\in\Omega\times[0,T].$
Finally, the value of the optimization problem admits the explicit expression
$V_{T,q}(\xi)={\mathbb{E}}[\xi]z^{(T,q)}_{0}+\dfrac{1}{2}\int_{0}^{T}{\mathrm{e}}^{-\rho s}(z^{(T,q)}_{s})^{2}\,{\mathrm{d}}s,$ (3.5)
and the optimally controlled state process $X^{\xi,\widehat{\mathbf{u}}^{(T,q)}}$ is given by
$X^{\xi,\widehat{\mathbf{u}}^{(T,q)}}_{s}=Y_{s}\xi+\int_{0}^{s}\dfrac{Y_{s}}{Y_{r}}z^{(T,q)}_{r}\,{\mathrm{d}}r,\quad s\in[0,T]\,,$ (3.6)
where $Y$ is the process defined in (2.5).
Proof.
Fix $\xi\in{\mathrm{L}}^{1}((\Omega,{\mathcal{F}}_{0},{\mathbb{P}});(0,+\infty))$. From (3.3) it immediately follows that
$V_{T,q}(\xi)={\mathbb{E}}[\xi]z^{(T,q)}_{0}+\sup_{{\mathbf{u}}\in{\mathcal{U}}_{T}}{\mathbb{E}}\left[\int_{0}^{T}{\mathrm{e}}^{-\rho s}\left\{z^{(T,q)}_{s}u_{s}-\dfrac{1}{2}u_{s}^{2}\right\}{\mathrm{d}}s\right].$ (3.7)
Hence, if we can find $\widehat{\mathbf{u}}^{(T,q)}\in{\mathcal{U}}_{T}$ that maximizes the integrand $z^{(T,q)}_{s}u_{s}-\dfrac{1}{2}u_{s}^{2}$, for ${\mathbb{P}}\otimes\mathrm{Leb}$-a.e. $(\omega,s)\in\Omega\times[0,T]$, then $\widehat{\mathbf{u}}^{(T,q)}$ it must be optimal. Clearly, the integrand is maximized in the sense above if we take $\widehat{\mathbf{u}}^{(T,q)}=z^{(T,q)}$. Its admissibility is a consequence of the fact that $z^{(T,q)}$ is bounded. Therefore, $\widehat{\mathbf{u}}^{(T,q)}$ is optimal, and clearly essentially unique in the sense specified above. Substituting its definition in (3.7) we immediately get (3.5) and (3.6). ∎
3.2. The infinite time horizon case
We analyze now the infinite time horizon case, i.e., we set $T=+\infty$. For each fixed $q\colon[0,+\infty)\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}((0,+\infty);(0,+\infty))$, we consider the problem of maximizing the functional
$J_{\infty,q}(\xi,{\mathbf{u}})\coloneqq{\mathbb{E}}\left[\int_{0}^{+\infty}{\mathrm{e}}^{-\rho s}\left\{X^{\xi,{\mathbf{u}}}_{s}q_{s}^{-\beta}-\dfrac{1}{2}u_{s}^{2}\right\}{\mathrm{d}}s\right],$ (3.8)
where $\xi$ satisfies Assumption 2.1, ${\mathbf{u}}$ is chosen in the class of admissible controls defined in (2), and $X^{\xi,{\mathbf{u}}}$ is the unique solution of SDE (2.1). We also introduce the value function corresponding to the optimization problem, namely,
$V_{\infty,q}(\xi)\coloneqq\sup_{{\mathbf{u}}\in{\mathcal{U}}_{\infty}}J_{\infty,q}(\xi,{\mathbf{u}}),\quad\xi\in{\mathrm{L}}^{1}((\Omega,{\mathcal{F}}_{0},{\mathbb{P}});(0,+\infty)).$ (3.9)
The next result is analogous to Proposition 3.1. The proof proceeds along the same lines and, thus, we omit it.
Proposition 3.4.
Fix a random variable $\xi$ satisfying Assumption 2.1 and ${\mathbf{u}}\in{\mathcal{U}}_{\infty}$. Consider a function $q\colon[0,+\infty)\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}_{\rho+\delta}((0,+\infty);(0,+\infty))$ and such that the function $z^{(\infty,q)}\colon[0,+\infty)\to(0,+\infty)$ given by
$z^{(\infty,q)}_{s}\coloneqq\int_{s}^{+\infty}{\mathrm{e}}^{-(\rho+\delta)(r-s)}q_{r}^{-\beta}\,{\mathrm{d}}r,\quad s\geq 0,$ (3.10)
is bounded on $[0,+\infty)$.
Then, the functional $J_{\infty,q}$ defined in (3.8) is finite and verifies
$J_{\infty,q}(\xi,{\mathbf{u}})={\mathbb{E}}[\xi]z^{(\infty,q)}_{0}+{\mathbb{E}}\left[\int_{0}^{+\infty}{\mathrm{e}}^{-\rho s}\left\{z^{(\infty,q)}_{s}u_{s}-\dfrac{1}{2}u_{s}^{2}\right\}{\mathrm{d}}s\right].$ (3.11)
Moreover, if $\xi,\xi^{\prime}$ both verify Assumption 2.1 and are such that ${\mathbb{E}}[\xi]={\mathbb{E}}[\xi^{\prime}]$, then $J_{\infty,q}(\xi,{\mathbf{u}})=J_{\infty,q}(\xi^{\prime},{\mathbf{u}})$, for any ${\mathbf{u}}\in{\mathcal{U}}_{\infty}$.
Remark 3.5.
Also in this case (cf. Remark 3.2) the last statement in Proposition 3.4 entails that the value function $V_{\infty,q}$ depends on the initial condition $\xi$ only through its average.
Thanks to Proposition 3.4, also in the infinite time horizon case we can find the optimal control for problem (3.9) and the explicit form of the corresponding value function, as stated in the next theorem. Its proof is omitted, being similar to that of Theorem 3.3.
Theorem 3.6.
Fix a random variable $\xi$ satisfying Assumption 2.1. Consider a function $q\colon[0,+\infty)\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}_{\rho+\delta}((0,+\infty);(0,+\infty))$ and such that the function $z^{(\infty,q)}$, defined in (3.10) is bounded on $[0,+\infty)$.
Then, $\widehat{\mathbf{u}}^{(\infty,q)}\coloneqq z^{(\infty,q)}\in{\mathcal{U}}_{\infty}$ is an optimal control for problem (3.9), which is deterministic and independent of $\xi$.
Moreover, $\widehat{\mathbf{u}}^{(\infty,q)}$ is essentially unique, i.e., if $\overline{\mathbf{u}}^{(\infty,q)}\in{\mathcal{U}}_{\infty}$ is an optimal control for problem (3.9) different from $\widehat{\mathbf{u}}^{(\infty,q)}$, then
$\overline{u}^{(\infty,q)}_{s}=\widehat{u}^{(\infty,q)}_{s},\quad\text{for ${\mathbb{P}}\otimes\mathrm{Leb}$-a.e. }(\omega,s)\in\Omega\times[0,+\infty).$
Finally, the value function of the optimization problem admits the explicit expression
$V_{\infty,q}(\xi)={\mathbb{E}}[\xi]z^{(\infty,q)}_{0}+\dfrac{1}{2}\int_{0}^{+\infty}{\mathrm{e}}^{-\rho s}(z^{(\infty,q)}_{s})^{2}\,{\mathrm{d}}s,$ (3.12)
and the optimally controlled state process $X^{\xi,\widehat{\mathbf{u}}^{(\infty,q)}}$ is given by
$X^{\xi,\widehat{\mathbf{u}}^{(\infty,q)}}_{s}=Y_{s}\xi+\int_{0}^{s}\dfrac{Y_{s}}{Y_{r}}z^{(\infty,q)}_{r}\,{\mathrm{d}}r,\quad s\geq 0,$ (3.13)
where $Y$ is the process defined in (2.5).
Remark 3.7.
All the results of this section can be straightforwardly extended to the case where the optimization problem starts at a time $t>0$. More precisely, setting the initial condition $X_{t}=x>0$ in SDE (2.1), we get that (2.6) becomes
${\mathbb{E}}[X^{\xi,{\mathbf{u}}}_{s}]={\mathbb{E}}[\xi]{\mathrm{e}}^{-\delta(s-t)}+{\mathbb{E}}\left[\int_{t}^{s}{\mathrm{e}}^{-\delta(s-r)}u_{r}\,{\mathrm{d}}r\right],\quad s\in[t,T],$
while the expressions of the functional $J_{T,q}$ and of the value function $V_{T,q}$ (now both dependent on $t$) given in (3.3) and (3.5), if $T<+\infty$, and in (3.11) and (3.12), if $T=+\infty$, become
$J_{T,q}(t,\xi,{\mathbf{u}})={\mathbb{E}}[\xi]{\mathrm{e}}^{-\rho t}z^{(T,q)}_{t}+{\mathbb{E}}\left[\int_{t}^{T}{\mathrm{e}}^{-\rho s}\left\{z^{(T,q)}_{s}u_{s}-\dfrac{1}{2}u_{s}^{2}\right\}{\mathrm{d}}s\right],$
$V_{T,q}(t,\xi)={\mathbb{E}}[\xi]\,{\mathrm{e}}^{-\rho t}z^{(T,q)}_{t}+\dfrac{1}{2}\int_{t}^{T}{\mathrm{e}}^{-\rho s}(z^{(T,q)}_{s})^{2}\,{\mathrm{d}}s,$
with optimal control $\widehat{\mathbf{u}}^{(T,q)}\coloneqq z^{(T,q)}$, where
$z^{(T,q)}_{s}\coloneqq\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta)(r-s)}q_{r}^{-\beta}\,{\mathrm{d}}r,\quad s\in[t,T].$
4. Existence and uniqueness of equilibria
In this section we discuss the existence and the uniqueness of an equilibrium for the mean-field game introduced in Section 2.
As usual in mean-field games theory, looking for an equilibrium boils down to finding a fixed point of a suitable map. According to Definition 2.3, in our case this map is $(q_{s})_{s\in[0,T]}\mapsto\left({\mathbb{E}}[X_{s}^{\xi,\widehat{\mathbf{u}}^{(T,q)}}]\right)_{s\in[0,T]}$, where $\widehat{\mathbf{u}}^{(T,q)}$ is the optimal control for the maximization problem studied in Section 3, whose expression is given either in Theorem 3.3, in the case $T<+\infty$, or in Theorem 3.6, in the case $T=+\infty$.
The next result formalizes this fact, showing also that the fixed point map is precisely the solution map of an integral equation.
Theorem 4.1.
Let us fix a random variable $\xi$ satisfying Assumption 2.1.
Consider the mean-field game introduced in Section 2 in the finite time horizon case, i.e., $T<+\infty$. Then,
  1. (i)
    If there exists an equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ in the sense of Definition 2.3, such that
    $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty)),$ (4.1)
    then $\widehat{q}$ is a solution to the integral equation
    $y_{s}={\mathrm{e}}^{-\delta s}{\mathbb{E}}[\xi]+\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}\int_{r}^{T}{\mathrm{e}}^{-(\rho+\delta)(u-r)}y_{u}^{-\beta}\,{\mathrm{d}}u\,{\mathrm{d}}r,\quad s\in[0,T]\,.$ (4.2)
  2. (ii)
    Vice versa, if there exist a unique solution $\widehat{q}$ to (4.2) satisfying (4.1), then there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})=(z^{(T,\widehat{q})},\widehat{q})$ of the mean-field game among all equilibria $(\widetilde{\mathbf{u}},\widetilde{q})$ such that $\widetilde{q}$ verifies (4.1), where $z^{(T,q)}$ is the function defined in (3.4).
Consider, instead, the same mean-field game in the infinite time horizon case, i.e., $T=+\infty$. Then,
  1. (iii)
    If there exists an equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ in the sense of Definition 2.3, such that
    $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}_{\rho+\delta}((0,+\infty);(0,+\infty))\quad\text{and}\quad z^{(\infty,\widehat{q})}\text{ is bounded on }[0,+\infty)\,,$ (4.3)
    where $z^{(\infty,q)}$ is the function defined in (3.10), then $\widehat{q}$ is a solution to the integral equation
    $y_{s}={\mathrm{e}}^{-\delta s}{\mathbb{E}}[\xi]+\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}\int_{r}^{+\infty}{\mathrm{e}}^{-(\rho+\delta)(u-r)}y_{u}^{-\beta}\,{\mathrm{d}}u\,{\mathrm{d}}r,\quad s\geq 0\,.$ (4.4)
  2. (iv)
    Vice versa, if there exist a unique solution $\widehat{q}$ to (4.4) satisfying (4.3), then there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})=(z^{(\infty,\widehat{q})},\widehat{q})$ of the mean-field game among all equilibria $(\widetilde{\mathbf{u}},\widetilde{q})$ such that $\widetilde{q}$ verifies (4.3).
Proof.
We prove only the first two statements in the finite time horizon case, as the case $T=+\infty$ is analogous.
(i). Let $(\widehat{\mathbf{u}},\widehat{q})$ be an equilibrium of the mean-field game, with $\widehat{\mathbf{u}}\in{\mathcal{U}}_{T}$ and $\widehat{q}\colon[0,T]\to(0,+\infty)$, such that $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$. Since point (i) of Definition 2.3 holds, from Theorem 3.3 we have that $\widehat{\mathbf{u}}=z^{(T,\widehat{q})}$, for ${\mathbb{P}}\otimes\mathrm{Leb}$-a.e. $(\omega,s)\in\Omega\times[0,T]$. Combining the expression of $z^{(T,q)}$, given in (3.4), and (2.6), we get
${\mathbb{E}}[X^{\xi,\widehat{\mathbf{u}}}_{s}]$ $={\mathbb{E}}[\xi]{\mathrm{e}}^{-\delta s}+{\mathbb{E}}\left[\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}\widehat{u}_{r}\,{\mathrm{d}}r\right]={\mathbb{E}}[\xi]{\mathrm{e}}^{-\delta s}+{\mathbb{E}}\left[\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}z^{(T,\widehat{q})}_{r}\,{\mathrm{d}}r\right]$
$={\mathrm{e}}^{-\delta s}{\mathbb{E}}[\xi]+\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}\int_{r}^{T}{\mathrm{e}}^{-(\rho+\delta)(u-r)}\widehat{q}_{u}^{-\beta}\,{\mathrm{d}}u\,{\mathrm{d}}r,\quad s\in[0,T].$ (4.5)
Therefore, by point (ii) of Definition 2.3, we get that $\widehat{q}$ is a solution to (4.2).
(ii). Let $\widehat{q}$ be the unique solution to (4.2) such that $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$. Define $\widehat{\mathbf{u}}\coloneqq z^{(T,\widehat{q})}$ according to (3.4). Then, by Theorem 3.3, we know that $\widehat{\mathbf{u}}$ is optimal for the problem of maximizing $J_{T,\widehat{q}}$, i.e., that point (i) of Definition 2.3 holds. We also have that equation (4.5) above is verified, and hence, since we assumed that $\widehat{q}$ solves (4.2), we get that $\widehat{q}_{s}={\mathbb{E}}[X^{\xi,\widehat{\mathbf{u}}}_{s}]$, for all $s\in[0,T]$, i.e., that point (ii) of Definition 2.3 also holds. This means that $(\widehat{\mathbf{u}},\widehat{q})$ is an equilibrium of the mean-field game.
If $(\widetilde{\mathbf{u}},\widetilde{q})$ is any other equilibrium of the mean-field game, such that $\widetilde{q}^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$, then necessarily $\widetilde{\mathbf{u}}=z^{(T,\widetilde{q})}$, for ${\mathbb{P}}\otimes\mathrm{Leb}$-a.e. $(\omega,s)\in\Omega\times[0,T]$, by Theorem 3.3. Therefore, we can repeat the same argument of the proof of point (i) to get that $\widetilde{q}$ is a solution to (4.2), and hence that $(\widetilde{\mathbf{u}},\widetilde{q})=(\widehat{\mathbf{u}},\widehat{q})$, since we assumed that (4.2) has a unique solution. ∎
Thanks to Theorem 4.1 we can reduce our search for a unique equilibrium of our mean-field game to proving existence and uniqueness of a solution to the integral equation (4.2), in the finite time horizon case, or (4.4), in the infinite time horizon case. This, in turn, is equivalent to showing existence and uniqueness of a solution $y\colon[0,T]\to(0,+\infty)$ to the integro-differential equation
$\begin{dcases}\dfrac{{\mathrm{d}}}{{\mathrm{d}}s}y_{s}=-\delta y_{s}+\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta)(u-s)}y_{u}^{-\beta}\,{\mathrm{d}}u,&s\in[0,T],\\ y_{0}=x\,.\end{dcases}$ (4.6)
for any initial datum $x>0$.
Indeed, this equivalence follows from the next two definitions and from Proposition 4.4 below, whose proof is omitted, being a standard result.
Definition 4.2.
A function $y\in{\mathrm{C}}^{1}([0,T];(0,+\infty))$ that satisfies (4.6) for all $s\in[0,T]$ is called a classical solution to (4.6).
Definition 4.3.
A function $y\in{\mathrm{C}}([0,T];(0,+\infty))$ that satifies the integral equation
$y_{s}={\mathrm{e}}^{-\delta s}x+\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}\int_{r}^{T}{\mathrm{e}}^{-(\rho+\delta)(u-r)}y_{u}^{-\beta}\,{\mathrm{d}}u\,{\mathrm{d}}r,\quad s\in[0,T],$
is called a mild solution to (4.6).
Proposition 4.4.
Any classical solution $y\in{\mathrm{C}}^{1}([0,T];(0,+\infty))$ to (4.6) is also a mild solution. Vice versa, any mild solution $y\in{\mathrm{C}}([0,T];(0,+\infty))$ to (4.6) is continuously differentiable in $[0,T]$ and also a classical solution.
Remark 4.5.
From now on, we will just speak of a solution (in either sense) to (4.6), unless otherwise specified.
It is convenient to introduce a further integro-differential equation, namely
$\begin{dcases}\dfrac{{\mathrm{d}}}{{\mathrm{d}}s}z_{s}={\mathrm{e}}^{(\rho+2\delta)s}\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}z_{u}^{-\beta}\,{\mathrm{d}}u,&s\in[0,T],\\ z_{0}=x\,,\end{dcases}$ (4.7)
and the corresponding integral equation
$z_{s}=x+\int_{0}^{s}{\mathrm{e}}^{(\rho+2\delta)r}\int_{r}^{T}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}z_{u}^{-\beta}\,{\mathrm{d}}u\,{\mathrm{d}}r,\quad s\in[0,T]\,,$ (4.8)
which we consider for any initial datum $x>0$.
The notions of classical and mild solution to (4.7) are analogous to those given in Definitions 4.2 and 4.3. Also in this case, a result similar to Proposition 4.4 shows that these two solution concepts are equivalent.
The following result, whose proof is omitted being based on routine computations, shows that solutions to (4.6) can be expressed in terms of those of (4.7) and vice versa. Thus, we can study either of these two equations, depending on convenience.
Proposition 4.6.
If $y\colon[0,T]\to(0,+\infty)$ is a solution to (4.6), then $z_{s}\coloneqq{\mathrm{e}}^{\delta s}y_{s}$, $s\in[0,T]$ is a solution to (4.7). Vice versa, if $z\colon[0,T]\to(0,+\infty)$ is a solution to (4.7), then $y_{s}\coloneqq{\mathrm{e}}^{-\delta s}z_{s}$, $s\in[0,T]$ is a solution to (4.6).
Remark 4.7.
Note that equations (4.6) and (4.7) do not fall into the standard theory of integro-differential equations, as we have an initial-value problem with the integral being backward in time.
Now, we proceed to study existence and uniqueness of a solution to (4.6) or, equivalently, (4.7). As in the previous section, we are going to divide our analysis into two subsections, the first devoted to the finite time horizon case, the second one to the infinite time horizon case.
4.1. The finite time horizon case
Let us fix $T<+\infty$ and $x>0$. We observe, first, that solutions of (4.7) satisfy suitable a priori estimates.
Proposition 4.8.
Let $z\in{\mathrm{C}}([0,T];(0,+\infty))$ be a solution to (4.7). Then, for all $s\in[0,T]$,
$z_{s}\geq\underline{\alpha}^{x}_{s},\qquad and\qquad z_{s}\leq\overline{\alpha}^{x}_{s},$ (4.9)
where
$\underline{\alpha}^{x}_{s}\coloneqq x,\quad s\in[0,T]\,,$ (4.10)
and, for any $s\in[0,T]$,
$\overline{\alpha}^{x}_{s}\coloneqq\begin{dcases}x-\dfrac{x^{-\beta}}{\rho+\delta-\delta\beta}\left\{\dfrac{1-{\mathrm{e}}^{\delta(1+\beta)s}}{\delta(1+\beta)}-{\mathrm{e}}^{-(\rho+\delta-\delta\beta)T}\dfrac{1-{\mathrm{e}}^{(\rho+2\delta)s}}{\rho+2\delta}\right\},&\text{if }\beta\neq 1+\dfrac{\rho}{\delta},\\ x-\dfrac{x^{-\beta}}{(\rho+2\delta)^{2}}\left\{[1+(\rho+2\delta)(T-s)]{\mathrm{e}}^{(\rho+2\delta)s}-[1+(\rho+2\delta)T]\right\},&\text{if }\beta=1+\dfrac{\rho}{\delta}.\end{dcases}$ (4.11)
Proof.
Since $z_{s}>0$, for all $s\in[0,T]$, the right hand side of (4.8) is positive, and hence, again for all $s\in[0,T]$,
$z_{s}\geq x,\qquad and\qquad z_{s}^{-\beta}\leq x^{-\beta},$ (4.12)
which provides us with the lower bound. To get the upper bound, it is enough to substitute the second inequality in (4.12) into the right hand side of (4.8) and compute the integrals. ∎
Given the result above our goal will be to show that (4.7) admits a unique solution in the (nonempty) set
${\mathcal{C}}_{x}\coloneqq\{f\in{\mathrm{C}}([0,T];(0,+\infty))\colon\underline{\alpha}^{x}_{s}\leq f_{s}\leq\overline{\alpha}^{x}_{s},\,\forall s\in[0,T]\}.$ (4.13)
Theorem 4.9.
For each fixed $x>0$, there exists a unique solution to the integro-differential equation (4.7) in the set ${\mathcal{C}}_{x}$ defined in (4.13), which is also twice continuously differentiable. Consequently, the integro-differential equation (4.6) admits a unique solution $y$. Such solution belongs to ${\mathrm{C}}^{2}([0,T];(0,+\infty))$ and satisfies
${\mathrm{e}}^{-\delta s}x\leq y_{s}\leq{\mathrm{e}}^{-\delta s}\overline{\alpha}^{x}_{s},\qquad s\in[0,T].$ (4.14)
Proof.
Fix $x>0$. Showing that (4.7) admits a unique solution is equivalent to proving that the operator $\Gamma$, defined on ${\mathrm{C}}([0,T];(0,+\infty))$ and given by
$\Gamma(f)_{s}\coloneqq x+\int_{0}^{s}{\mathrm{e}}^{(\rho+2\delta)r}\int_{r}^{T}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}f_{u}^{-\beta}\,{\mathrm{d}}u\,{\mathrm{d}}r,\quad s\in[0,T],$ (4.15)
admits a unique fixed point in ${\mathcal{C}}_{x}$.
It is easy to verify that ${\mathcal{C}}_{x}$ is a closed, bounded, and convex subset of ${\mathrm{C}}([0,T];(0,+\infty))$ (which is endowed with the usual sup-norm). Moreover, $\Gamma$ maps ${\mathcal{C}}_{x}$ into itself. Indeed, from (4.15) we readily verify that, for all $f\in{\mathcal{C}}_{x}$ and all $s\in[0,T]$, $\Gamma(f)_{s}\geq\underline{\alpha}^{x}_{s}$, since $f$ is a positive function; moreover, noting that for all $f,g\in{\mathcal{C}}_{x}$
$f_{s}\leq g_{s}\quad\Longrightarrow\quad\Gamma(f)_{s}\geq\Gamma(g)_{s},\quad s\in[0,T],$
we immediately check that $\Gamma(f)_{s}\leq\Gamma(\underline{\alpha}^{x})_{s}=\overline{\alpha}^{x}_{s}$. Finally, $\Gamma$ is compact in ${\mathcal{C}}_{x}$ as a consequence of the Ascoli-Arzelà theorem. Indeed, any sequence in ${\mathcal{C}}_{x}$ is equi-bounded, thanks to the definition of ${\mathcal{C}}_{x}$, and equi-continuous since, for all $f\in{\mathcal{C}}_{x}$, $\Gamma(f)$ is differentiable and
$\sup_{f\in{\mathcal{C}}}\sup_{s\in[0,T]}\lvert\Gamma(f)_{s}^{\prime}\rvert=\sup_{f\in{\mathcal{C}}}\sup_{s\in[0,T]}\left\{{\mathrm{e}}^{(\rho+2\delta)s}\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}f_{u}^{-\beta}\,{\mathrm{d}}u\right\}\leq\sup_{s\in[0,T]}\frac{{\mathrm{d}}}{{\mathrm{d}}s}\overline{\alpha}^{x}_{s}<+\infty.$
Therefore, by Schauder’s fixed point theorem the map $\Gamma$ has a fixed point $z\in{\mathcal{C}}_{x}$, i.e., $z$ verifies (4.8). Hence, $z$ is a solution of equation (4.7). The fact that $z\in{\mathrm{C}}^{2}([0,T];(0,+\infty))$ simply follows by differentiating twice (4.8).
To show uniqueness, assume that two different solutions $z,\widetilde{z}\in{\mathcal{C}}_{x}$ of (4.7) exist. It is convenient to write the integro-differential equation satisfied by $w_{s}\coloneqq z_{T-s}$ and $\widetilde{w}_{s}\coloneqq\widetilde{z}_{T-s}$, $s\in[0,T]$,
$\begin{dcases}\dfrac{{\mathrm{d}}}{{\mathrm{d}}s}w_{s}=-{\mathrm{e}}^{\delta(1+\beta)(T-s)}\int_{0}^{s}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)(s-r)}w_{r}^{-\beta}\,{\mathrm{d}}r,&s\in[0,T],\\ w_{T}=x\,.\end{dcases}$ (4.16)
From this equation we deduce that
$w^{\prime}_{s}-\widetilde{w}^{\prime}_{s}=-{\mathrm{e}}^{\delta(1+\beta)(T-s)}\int_{0}^{s}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)(s-r)}\left[w_{r}^{-\beta}-\widetilde{w}_{r}^{-\beta}\right]{\mathrm{d}}r,\quad s\in[0,T]\,.$ (4.17)
The following two cases may occur:
Case 1. Consider $w_{0}\neq\widetilde{w}_{0}$ and assume, without loss of generality, that $w_{0}>\widetilde{w}_{0}$. Let us define
$s^{\star}\coloneqq\inf\{u>0\colon w_{u}=\widetilde{w}_{u}\}\,.$
For each fixed $s\in[0,s^{\star})$, we have that $w_{s}>\widetilde{w}_{s}$ and, hence, $w_{r}^{-\beta}-\widetilde{w}_{r}^{-\beta}<0$, for all $r\in[0,s]$; thus, we deduce from (4.17) that $w^{\prime}_{s}-\widetilde{w}^{\prime}_{s}>0$, i.e., that the distance between the two solutions increases in time. Therefore, $s^{\star}=+\infty$, but this contradicts the fact that $w_{T}=\widetilde{w}_{T}$.
Case 2. Consider $w_{0}=\widetilde{w}_{0}$ and note that both $w$ and $\widetilde{w}$ are bounded below by the constant $x$, since $z,\widetilde{z}\in{\mathcal{C}}_{x}$. Observe, also, that the function $w\mapsto w^{-\beta}$ is Lipschitz in $[x,+\infty)$, with Lipschitz constant $L\coloneqq x^{-\beta-1}$. Therefore, integrating (4.17) on $[0,s]$ we get
$w_{s}-\widetilde{w}_{s}=-\int_{0}^{s}{\mathrm{e}}^{\delta(1+\beta)(T-u)}\int_{0}^{u}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)(u-r)}\left[w_{r}^{-\beta}-\widetilde{w}_{r}^{-\beta}\right]{\mathrm{d}}r\,{\mathrm{d}}u,\quad s\in[0,T]\,,$
whence
$\lvert w_{s}-\widetilde{w}_{s}\rvert\leq L\int_{0}^{s}\left\{{\mathrm{e}}^{\delta(1+\beta)(T-u)}\int_{0}^{u}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)(u-r)}\,{\mathrm{d}}r\right\}\sup_{r\in[0,u]}\lvert w_{r}-\widetilde{w}_{r}\rvert\,{\mathrm{d}}u,\quad s\in[0,T]\,.$
Since $s\mapsto\sup_{r\in[0,s]}\lvert w_{r}-\widetilde{w}_{r}\rvert$ is bounded on $[0,T]$, applying Gronwall’s lemma we get that $w_{s}=\widetilde{w}_{s}$ for all $s\in[0,T]$, which contradicts the assumption that $w$ and $\widetilde{w}$ are different.
Therefore, equation (4.7) admits a unique solution in ${\mathcal{C}}_{x}$. Finally, from Proposition 4.6, we immediately deduce that also (4.6) has a unique solution in ${\mathrm{C}}^{2}([0,T];(0+\infty))$ satisfying the bounds in (4.14). ∎
Thanks to the result above, we can establish the following.
Proposition 4.10.
For each fixed random variable $\xi$ satisfying Assumption 2.1, there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ for the mean-field game with finite time horizon, among all equilibria verifying the integrability condition (4.1).
More precisely, $\widehat{\mathbf{u}}=z^{(T,\widehat{q})}$, where $z^{(T,\widehat{q})}$ is the function defined in (3.4), and $\widehat{q}$ is the unique solution to (4.6), with initial condition $x={\mathbb{E}}[\xi]$.
Proof.
Let $\widehat{q}$ be the unique solution to (4.6), with $x={\mathbb{E}}[\xi]$. Since $\widehat{q}$ satisfies the bounds (4.14) and since the functions $s\mapsto{\mathrm{e}}^{-\delta s}x$ and $s\mapsto{\mathrm{e}}^{-\delta s}\overline{\alpha}_{s}^{x}$ are continuous on $[0,T]$, hence therein bounded, we deduce that also $\widehat{q}$ is bounded on $[0,T]$. More precisely,
$0<m\leq\widehat{q}_{s}\leq M,\quad\forall s\in[0,T],\quad\text{with }m\coloneqq{\mathrm{e}}^{-\delta T}x\,\text{ and }M\coloneqq\sup_{s\in[0,T]}{\mathrm{e}}^{-\delta s}\overline{\alpha}_{s}^{x}\,.$ (4.18)
Therefore, $\widehat{q}$ satisfies the integrability condition (4.1) and applying Theorem 4.1-(ii) we get the result. ∎
4.2. The infinite time horizon case
We discuss now the infinite time horizon case, i.e., we fix $T=+\infty$. Here we work under the assumption $\beta<1+\frac{\rho}{\delta}$, which will be used to guarantee convergence of the integral appearing in (4.7).
A reasoning similar to the one used to prove Proposition 4.8 shows that the following holds.
Proposition 4.11.
Let $z\in{\mathrm{C}}([0,+\infty);(0,+\infty))$ be a solution to (4.7), with $T=+\infty$, and assume that $\beta<1+\frac{\rho}{\delta}$. Then, for all $s\geq 0$,
$z_{s}\geq\underline{\alpha}^{x,\infty}_{s},\qquad and\qquad z_{s}\leq\overline{\alpha}^{x,\infty}_{s}.$ (4.19)
where
$\underline{\alpha}^{x,\infty}_{s}\coloneqq x,\quad s\geq 0\,,$ (4.20)
and
$\overline{\alpha}^{x,\infty}_{s}\coloneqq x-\dfrac{x^{-\beta}}{\rho+\delta-\delta\beta}\left\{\dfrac{1-{\mathrm{e}}^{\delta(1+\beta)s}}{\delta(1+\beta)}\right\},\quad s\geq 0.$ (4.21)
We aim at showing that (4.7) admits a unique solution in the (nonempty) set
${\mathcal{C}}_{x,\infty}\coloneqq\{f\in{\mathrm{C}}([0,+\infty);(0,+\infty))\colon\underline{\alpha}^{x,\infty}_{s}\leq f_{s}\leq\overline{\alpha}^{x,\infty}_{s},\,\forall s\geq 0\}.$ (4.22)
Note, in particular, that for any $f\in{\mathcal{C}}_{x,\infty}$ we have that
$\int_{s}^{+\infty}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}f_{u}^{-\beta}\,{\mathrm{d}}u\leq\int_{s}^{+\infty}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}x^{-\beta}\,{\mathrm{d}}u=\frac{x^{-\beta}}{\rho+\delta-\delta\beta}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)s}<+\infty,\quad\forall s\geq 0,$
as $f$ is bounded below by the constant $x$ and we assumed $\beta<1+\frac{\rho}{\delta}$.
Theorem 4.12.
Let $T=+\infty$ and assume that $\beta<1+\frac{\rho}{\delta}$. Then, for each fixed $x>0$, there exists a unique solution to the integro-differential equation (4.7) in the set ${\mathcal{C}}_{x,\infty}$ defined in (4.22), which is also twice continuously differentiable. Moreover, the integro-differential equation (4.6) admits a unique solution $y$. Such solution belongs to ${\mathrm{C}}^{2}([0,+\infty);(0,+\infty))$ and satisfies
${\mathrm{e}}^{-\delta s}x\leq y_{s}\leq{\mathrm{e}}^{-\delta s}\left[x-\dfrac{x^{-\beta}}{\delta(1+\beta)(\rho+\delta-\delta\beta)}\right]+\dfrac{{\mathrm{e}}^{\delta\beta s}x^{-\beta}}{\delta(1+\beta)(\rho+\delta-\delta\beta)},\quad s\geq 0.$ (4.23)
Finally, if $x=y_{\infty}\coloneqq(\delta(\rho+\delta))^{-\frac{1}{\beta+1}}$, then the function $y_{s}=y_{\infty}$, for all $s\geq 0$, is the unique constant solution to (4.6). If, instead, $x\neq y_{\infty}$, then the corresponding solution $y$ to (4.6) is strictly monotone and such that
$\lim_{s\to+\infty}y_{s}=y_{\infty}\,.$
Proof.
We divide the proof as follows. First, we show existence and uniqueness of a solution to (4.7) in the set ${\mathcal{C}}_{x,\infty}$, for each fixed $x>0$. As in the finite time horizon case, the fact that such solution is twice continuously differentiable follows by differentiation of (4.8). Then, it immediately follows from Proposition 4.6 that (4.6) admits a unique solution $y\in{\mathrm{C}}^{2}([0,+\infty);(0,+\infty))$ satisfying the bounds given in (4.23). Finally, we will show that the constant function equal to $y_{\infty}$ is the unique constant solution to (4.6), and that the solutions starting at $x\neq y_{\infty}$ are monotone in time and converge to $y_{\infty}$.
Existence. Fix $x>0$, and consider the family of solutions ${\mathcal{Z}}\coloneqq\{z^{(n)}\}_{n\geq 1}$, where $z^{(n)}$ is the solution to the finite horizon problems with terminal time $T=n$, i.e.,
$\begin{dcases}\dfrac{{\mathrm{d}}}{{\mathrm{d}}s}z^{(n)}_{s}={\mathrm{e}}^{(\rho+2\delta)s}\int_{s}^{n}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}\bigl{(}z_{u}^{(n)}\bigr{)}^{-\beta}\,{\mathrm{d}}u,&s\in[0,n],\\ z_{0}=x\,.\end{dcases}$ (4.24)
We extend each of them by continuity taking them to be constant over $(n,+\infty)$, so that $z^{(n)}\in{\mathcal{C}}_{x,\infty}$, as $\overline{\alpha}_{s}^{x}\leq\overline{\alpha}_{s}^{x,\infty}$, for all $s\in[0,n]$. Differently from the finite horizon case, here we cannot directly apply the Ascoli-Arzelà theorem, as the infinite horizon does not allow to get the same equicontinuity estimates as in the proof of Theorem 4.9. Hence, use a weaker compactness criterion based on the Fréchet-Kolmogorov theorem.
Let us define the set
${\mathrm{L}}^{1}_{\rho+2\delta}([0,\infty);\mathbb{R})\coloneqq\left\{f\colon[0,\infty)\to\mathbb{R}\,\colon\,\int_{0}^{\infty}{\mathrm{e}}^{-(\rho+2\delta)s}\lvert f_{s}\rvert\,{\mathrm{d}}s<+\infty\right\}\,,$
endowed with the usual weighted norm $\lVert\cdot\rVert_{1,\rho+2\delta}$. Clearly ${\mathcal{Z}}\subset{\mathrm{L}}^{1}_{\rho+2\delta}([0,\infty);\mathbb{R})$.
We observe that, given the bounds in (4.22), the family ${\mathcal{Z}}$ is equibounded. We want to show that it is also equi-integrable in ${\mathrm{L}}^{1}_{\rho+2\delta}$. For any $s\geq 0$ and $h>0$ we have that
$z^{(n)}_{s+h}-z^{(n)}_{s}=\mathbf{1}_{[0,n-h]}(s)\int_{s}^{s+h}\bigl{(}z^{(n)}_{r}\bigr{)}^{\prime}\,{\mathrm{d}}r+\mathbf{1}_{[n-h,n]}(s)\int_{s}^{n}\bigl{(}z^{(n)}_{r}\bigr{)}^{\prime}\,{\mathrm{d}}r.$
Let $K\coloneqq\frac{x^{-\beta}}{\rho+\delta-\delta\beta}$. Using the fact that $0<\bigl{(}z^{(n)}_{s}\bigr{)}^{\prime}\leq K{\mathrm{e}}^{\delta(1+\beta)s}$, for all $s\geq 0$, and all $n\geq 1$, we get that, for any $h>0$ small enough,
$\phantom{\leq}\lVert z^{(n)}_{\cdot+h}-z^{(n)}_{\cdot}\rVert_{1,\rho+2\delta}=\int_{0}^{+\infty}{\mathrm{e}}^{-(\rho+2\delta)s}\lvert z^{(n)}_{s+h}-z^{(n)}_{s}\rvert\,{\mathrm{d}}s$
$\leq\int_{0}^{n-h}{\mathrm{e}}^{-(\rho+2\delta)s}\int_{s}^{s+h}\bigl{(}z^{(n)}_{r}\bigr{)}^{\prime}\,{\mathrm{d}}r\,{\mathrm{d}}s+\int_{n-h}^{n}{\mathrm{e}}^{-(\rho+2\delta)s}\int_{s}^{n}\bigl{(}z^{(n)}_{r}\bigr{)}^{\prime}\,{\mathrm{d}}r$
$\leq\dfrac{K({\mathrm{e}}^{\delta(1+\beta)h}-1)}{\delta(1+\beta)}\int_{0}^{n-h}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)s}\,{\mathrm{d}}s$
$\qquad+\dfrac{K{\mathrm{e}}^{\delta(1+\beta)n}}{\delta(1+\beta)}\int_{n-h}^{n}{\mathrm{e}}^{-(\rho+2\delta)s}\,{\mathrm{d}}s\underbrace{-\dfrac{K}{\delta(1+\beta)}\int_{n-h}^{n}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)s}\,{\mathrm{d}}s}_{\leq 0}$
$\leq\dfrac{K({\mathrm{e}}^{\delta(1+\beta)h}-1)}{\delta(1+\beta)(\rho+\delta-\delta\beta)}[1\underbrace{-{\mathrm{e}}^{-(\rho+\delta-\delta\beta)(n-h)}}_{\leq 0}]+\dfrac{K{\mathrm{e}}^{-(\rho+\delta-\delta\beta)n}}{\delta(1+\beta)(\rho+2\delta)}[{\mathrm{e}}^{(\rho+2\delta)h}-1]$
$\leq\dfrac{K({\mathrm{e}}^{\delta(1+\beta)h}-1)}{\delta(1+\beta)(\rho+\delta-\delta\beta)}+\dfrac{K}{\delta(1+\beta)(\rho+2\delta)}[{\mathrm{e}}^{(\rho+2\delta)h}-1],\quad\forall n\geq 1\,.$
Therefore, we obtain that
$\lim_{h\to 0^{+}}\sup_{n\geq 1}\lVert z^{(n)}_{\cdot+h}-z^{(n)}_{\cdot}\rVert_{1,\rho+2\delta}=0,$
i.e., that the family ${\mathcal{Z}}$ is equi-integrable in ${\mathrm{L}}^{1}_{\rho+2\delta}$.
By the Frechet-Kolmogorov theorem, it follows that ${\mathcal{Z}}$ is relatively compact, so we can extract a subsequence $\{z^{(n)}\}_{n\geq 1}\subset{\mathcal{Z}}$, still labeled (with an abuse of notation) by $\{z^{(n)}\}_{n\geq 1}$, such that $z^{(n)}\longrightarrow z^{(\infty)}$ with respect to $\lVert\cdot\rVert_{1,\rho+2\delta}$, as $n\to\infty$. Then, we can extract a sub-subsequence, still labeled (with an abuse of notation) by $\{z^{(n)}\}_{n\geq 1}$, such that $z^{(n)}_{s}\longrightarrow z^{(\infty)}_{s}$, for almost every $s\geq 0$, as $n\to\infty$. Clearly, the limit $z^{(\infty)}$ belongs to ${\mathrm{L}}^{1}_{\rho+2\delta}$ and, since $z^{(n)}\in{\mathcal{C}}_{x,\infty}$, for any $n\geq 1$, we have that $\underline{\alpha}_{s}^{x,\infty}\leq z^{(\infty)}_{s}\leq\overline{\alpha}_{s}^{x,\infty}$, for almost every $s\geq 0$, where $\underline{\alpha}^{x,\infty}$ and $\overline{\alpha}^{x,\infty}$ are the functions defined in (4.20) and (4.21). Now we use (4.7) to get that also the derivative $\{\left(z^{(n)}\right)^{\prime}\}$ of the last subsequence $\{z^{(n)}\}$ converges a.e. to the function
$w^{(\infty)}_{s}\coloneqq{\mathrm{e}}^{(\rho+2\delta)s}\int_{s}^{+\infty}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}\left(z^{(\infty)}_{u}\right)^{-\beta}\,{\mathrm{d}}u,\quad s\geq 0\,,$
which belongs to ${\mathrm{L}}^{1}_{\rho+2\delta}$. By standard computations, it follows that $\{\left(z^{(n)}\right)^{\prime}\}\subset{\mathrm{L}}^{1}_{\rho+2\delta}$ and that
$0<\left(z^{(n)}_{s}\right)^{\prime}\leq\dfrac{x^{-\beta}}{\rho+\delta-\delta\beta}{\mathrm{e}}^{\delta(1+\beta)s}\eqqcolon g_{s},\quad\forall s\geq 0,\,\forall n\geq 1,$
with $g\in{\mathrm{L}}^{1}_{\rho+2\delta}$. Therefore, the sequence of derivatives $\{\left(z^{(n)}\right)^{\prime}\}$ converges in ${\mathrm{L}}^{1}_{\rho+2\delta}$ to $w^{(\infty)}$. We deduce that the subsequence $\{z^{(n)}\}$ converges in the weighted Sobolev space $W^{1,1}_{\rho+2\delta}$ (cf. [4, Chapter 8, Remark 4]). By completeness of this space, we get that $w^{(\infty)}_{s}=(z^{(\infty)}_{s})^{\prime}$, for a.e. $s\geq 0$. Moreover, using the same argument as in the proof of [4, Theorem 8.2], we get that there exists a function $\overline{z}^{(\infty)}\in{\mathrm{C}}([0,+\infty);(0,+\infty))$ such that $\overline{z}^{(\infty)}_{s}=z^{(\infty)}_{s}$, for almost every $s\geq 0$, and
$\overline{z}^{(\infty)}_{s}=x+\int_{0}^{s}(z^{(\infty)}_{u})^{\prime}\,{\mathrm{d}}u=x+\int_{0}^{s}w^{(\infty)}_{u}\,{\mathrm{d}}u,\quad\forall s\geq 0\,.$
Substituting the expression of $w^{(\infty)}$ in the previous equality we get that
$\overline{z}^{(\infty)}_{s}=x+\int_{0}^{s}{\mathrm{e}}^{(\rho+2\delta)r}\int_{r}^{+\infty}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}\left(\overline{z}^{(\infty)}_{u}\right)^{-\beta}\,{\mathrm{d}}u\,{\mathrm{d}}r\quad s\geq 0,$
that is, $\overline{z}^{(\infty)}\in{\mathcal{C}}_{x,\infty}$ and solves (4.7).
Uniqueness. As we observed at the beginning of the proof, any solution $z$ to (4.7) is (at least) twice differentiable. Differentiating (4.7) with respect to the time variable and denoting by and ′′ first- and second-order derivatives with respect to this variable, we have that $z$ verifies
$\begin{dcases}z^{\prime\prime}_{s}=(\rho+2\delta)z^{\prime}_{s}-{\mathrm{e}}^{\delta(1+\beta)s}z_{s}^{-\beta},&s\geq 0,\\ z_{0}=x\,.\end{dcases}$ (4.25)
The equation above is a second-order ODE with locally Lipschitz coefficients over $(s,z)\in[0,+\infty)\times(0,+\infty)$, to which we can associate the following initial value problem
$\begin{dcases}z^{\prime\prime}_{s}=(\rho+2\delta)z^{\prime}_{s}-{\mathrm{e}}^{\delta(1+\beta)s}z_{s}^{-\beta},&s\geq 0,\\ z_{0}=x\,,\\ z^{\prime}_{0}=\zeta>0\,.\end{dcases}$ (4.26)
Note that we only need to consider $\zeta>0$, since solutions to (4.7) are such that their first derivative is strictly positive for all $s\geq 0$. For each fixed $x>0$ and $\zeta>0$, (4.26) has a unique solution $z^{x,\zeta}$ on the maximal interval $[0,\tau^{*}(x,\zeta))$, with $\tau^{*}(x,\zeta)\leq+\infty$.
Since any solution $z$ to (4.7) also verifies (4.25), it also satisfies (4.26) for some $\overline{\zeta}>0$. Therefore, if we show that for each given $x>0$, there exists a unique $\zeta>0$ such that (4.26) has a unique global solution $z^{x,\zeta}\in\mathcal{C}_{x,\infty}$, then necessarily $\zeta=\overline{\zeta}$ and $z^{x,\zeta}$ must also be the unique solution to (4.7). The idea is, thus, to study, for each fixed $x>0$, the dependence of the solution $z^{x,\zeta}$ to (4.26) on $\zeta>0$.
Let us fix $x>0$. Standard results (see, e.g., [11, Theorem 7,5, Chapter 1]) ensure that the solution $z^{x,\zeta}$ to (4.26) and its first- and second-order derivatives with respect to the time variable are differentiable with respect to $\zeta$ and that the order of differentiation can be exchanged.
Therefore, defining $\widehat{z}_{s}({\zeta})\coloneqq\frac{\partial}{\partial\zeta}z_{s}^{x,\zeta}$, for all $s\geq 0$, we get from (4.26) that $\widehat{z}(\zeta)$ solves the initial value problem
$\begin{dcases}\widehat{z}^{\prime\prime}_{s}(\zeta)=(\rho+2\delta)\widehat{z}^{\prime}_{s}(\zeta)+\beta{\mathrm{e}}^{\delta(1+\beta)s}z_{s}(\zeta)^{-\beta-1}\widehat{z}_{s}(\zeta),&s\geq 0,\\ \widehat{z}_{0}(\zeta)=0\,,\\ \widehat{z}^{\prime}_{0}(\zeta)=1\,,\end{dcases}$ (4.27)
where $z_{s}(\zeta)\coloneqq z_{s}^{x,\zeta}$. Let
$s^{*}(\zeta)\coloneqq\inf\{s\geq 0\colon\widehat{z}^{\prime}_{s}(\zeta)\leq 0\}>0.$
By definition $z^{\prime}_{s}(\zeta)>0$, for all $s\in[0,s^{*}(\zeta))$, which implies
$\widehat{z}_{s}(\zeta)>0,\qquad\forall s\in[0,s^{*}(\zeta)).$
Therefore, using (4.27) we get that $\widehat{z}^{\prime\prime}_{s}(\zeta)>0$, for all $s\in[0,s^{*}(\zeta))$. It follows that
$\widehat{z}^{\prime}_{s}(\zeta)\geq 1,\qquad\forall s\in[0,s^{*}(\zeta)),$
which implies that $s^{*}(\zeta)=+\infty$ and, in turn, that $\widehat{z}_{s}(\zeta)\geq s$, for all $s\geq 0$.
Using again (4.27), we then get
$\begin{dcases}\widehat{z}^{\prime\prime}_{s}(\zeta)\geq(\rho+2\delta)\widehat{z}^{\prime}_{s}(\zeta),&s\geq 0,\\ \widehat{z}^{\prime}_{0}(\zeta)=1\,,\end{dcases}$ (4.28)
which, implies, by comparison,
$\widehat{z}^{\prime}_{s}(\zeta)\geq{\mathrm{e}}^{(\rho+2\delta)s},\qquad\forall s\geq 0,$
and, consequently,
$\widehat{z}_{s}(\zeta)\geq\frac{1}{\rho+2\delta}[{\mathrm{e}}^{(\rho+2\delta)s}-1],\qquad\forall s\geq 0.$
Thus, for every $a<b$, we have
$z_{s}(b)-z_{s}(a)\geq\int_{a}^{b}\hat{z}_{s}(\zeta)\,{\mathrm{d}}\zeta\geq(b-a)\frac{1}{\rho+2\delta}[{\mathrm{e}}^{(\rho+2\delta)s}-1],\qquad\forall s\geq 0.$
We know that, for some $\overline{\zeta}>0$, there exists a global solution $z(\overline{\zeta})\in\mathcal{C}_{x,\infty}$ to (4.26), since any solution $z$ to (4.7) also verifies this equation. Using the estimate above and (4.22), we get, for all $b>a\coloneqq\overline{\zeta}$,
$z_{s}(b)\geq z_{s}(\overline{\zeta})+(b-\overline{\zeta})[{\mathrm{e}}^{(\rho+2\delta)s}-1]\geq x+(b-\overline{\zeta})\frac{1}{\rho+2\delta}[{\mathrm{e}}^{(\rho+2\delta)s}-1],\qquad\forall s\geq 0.$
This implies that $z(b)$ would grow exponentially at a rate at least $\rho+2\delta>\delta(1+\beta)$, which contradicts the fact that solutions in (4.22) need to grow at a rate at most $\delta(1+\beta)$.
Similarly, we get, for all $a<\overline{\zeta}\eqqcolon b$,
$z_{s}(a)\leq z_{s}(\overline{\zeta})-(\overline{\zeta}-a)[{\mathrm{e}}^{(\rho+2\delta)s}-1]\leq\overline{\alpha}_{s}^{x,\infty}-(\overline{\zeta}-a)\frac{1}{\rho+2\delta}[{\mathrm{e}}^{(\rho+2\delta)s}-1],\qquad\forall s\geq 0.$
This implies that $z(a)$ becomes negative in finite time, which contradicts the fact that solutions in (4.22) are positive.
By arbitrariness of $b>\overline{\zeta}$ and $a<\overline{\zeta}$, we conclude.
Convergence. As recalled at the beginning of the proof, uniqueness of the solution to (4.7) implies uniqueness of the solution to (4.6), which is also twice continuously differentiable. Differentiating (4.6), we get the second-order ODE
$y^{\prime\prime}_{s}=\rho y^{\prime}_{s}+(\rho+\delta)\delta y_{s}-y_{s}^{-\beta},\quad s\geq 0.$ (4.29)
Hence, any solution to (4.6) is also a solution to (4.29). We observe that $y_{\infty}$ is the unique solution to the algebraic equation
$(\rho+\delta)\delta y-y^{-\beta}=0.$ (4.30)
Hence, the function $y_{s}=y_{\infty}$, for all $s\geq 0$, is the unique stationary solution to (4.6).
Now, we show monotonicity and convergence of the solution to (4.6) when $x\neq y_{\infty}$. We prove the case $x>y_{\infty}$, as the other one can be established by similar arguments.
Let $y$ be the unique solution to (4.6) and fix $x>y_{\infty}$. We prove, first, that $y_{s}>y_{\infty}$, for all $s\geq 0$. Let us define
$s_{0}\coloneqq\inf\{s\geq 0\colon y_{s}=y_{\infty}\}\,,$
and assume, by contradiction, that $s_{0}<+\infty$. We have two cases:
  • (i)
    $y_{s_{0}}^{\prime}<0$;
  • (ii)
    $y_{s_{0}}^{\prime}=0$.
Consider the first case. By the flow property, the function $\widetilde{y}_{s}\coloneqq y_{s-s_{0}}$, $s\geq s_{0}$, solves (4.6) with initial condition equal to $y_{\infty}$ and it satisfies the a priori bounds given in (4.23) (with $x=y_{\infty}$). However, this solution is different from the constant one, which is the unique solution to (4.6) with initial condition $y_{\infty}$ and satisfying the aforementioned bounds. Hence we have a contradiction.
The second case leads to a contradiction as well. Indeed, we would have two different solutions in the interval $[0,s_{0}]$ to the Cauchy problem (4.29), with terminal conditions $y_{s_{0}}=y_{\infty}$ and $y^{\prime}_{s_{0}}=0$, which clearly admits unique solution, that is, $y_{s}=y_{\infty}$, for all $s\geq 0$.
Hence, we have proved that $y_{s}>y_{\infty}$, for all $s\geq 0$. This implies, by strict monotonicity of the map $y\mapsto(\rho+\delta)\delta y-y^{-\beta}$ and (4.30), that
$(\rho+\delta)\delta y_{s}-y_{s}^{-\beta}>(\rho+\delta)\delta y_{\infty}-y_{\infty}^{-\beta}=0,\quad\forall s\geq 0\,,$ (4.31)
and hence, from (4.29), that
$y_{s}^{\prime\prime}>\rho y^{\prime}_{s},\quad\forall s\geq 0.$ (4.32)
Now we show that $y^{\prime}_{s}<0$, for all $s\geq 0$. Let
$s_{1}\coloneqq\inf\{s\geq 0\colon y^{\prime}_{s}\geq 0\}.$
Assume, by contradiction, that $s_{1}<+\infty$. Then, from (4.32) we get that $y^{\prime\prime}_{s_{1}}>0$. It follows that there exist $\eta,\varepsilon>0$ such that $y^{\prime}_{s_{1}+\varepsilon}=\eta>0$. Using (4.32) again, we get
$\begin{dcases}y^{\prime\prime}_{s}>\rho y^{\prime}_{s},&\forall s\geq s_{1}+\varepsilon,\\ y^{\prime}_{s_{1}+\varepsilon}=\eta>0.\end{dcases}$ (4.33)
Therefore, solving the corresponding Cauchy problem and using the comparison, we get
$y^{\prime}_{s}\geq\eta{\mathrm{e}}^{\rho[s-(s_{1}+\varepsilon)]}>0,\quad s\geq s_{1}+\varepsilon.$ (4.34)
Next, differentiating (4.29), we get that $y\in{\mathrm{C}}^{3}([0,+\infty);(0,+\infty))$ and that it satisfies
$y^{\prime\prime\prime}_{s}=\rho y^{\prime\prime}_{s}+(\rho+\delta)\delta y^{\prime}_{s}+\beta y_{s}^{-\beta-1}y^{\prime}_{s},\quad s\geq 0.$ (4.35)
Since, by (4.34), we know that $y_{s}^{\prime}>0$ for all $s\geq s_{1}+\varepsilon$, we deduce that
$\begin{dcases}y_{s}^{\prime\prime\prime}>\rho y_{s}^{\prime\prime}+(\rho+\delta)\delta y^{\prime}_{s},&\forall s\geq s_{1}+\varepsilon,\\ y^{\prime}_{s_{1}+\varepsilon}=\eta,\\ y^{\prime\prime}_{s_{1}+\varepsilon}=\kappa,\end{dcases}$
with $\eta>0$ and $\kappa>\eta\rho>0$, by (4.33). Solving the corresponding Cauchy problem and using the comparison, we get
$y^{\prime}_{s}>\dfrac{\eta\delta+\kappa}{\rho+2\delta}{\mathrm{e}}^{(\rho+\delta)[s-(s_{1}+\varepsilon)]}+\dfrac{\eta(\rho+\delta)-\kappa}{\rho+2\delta}{\mathrm{e}}^{-\delta[s-(s_{1}+\varepsilon)]},\quad\forall s\geq s_{1}+\varepsilon\,.$
This implies that $y$ grows, in the long run, at least with rate $\rho+\delta$ which is strictly bigger than $\delta\beta$, as prescribed by the admissibility condition in (4.23). The contradiction follows, and hence we proved that $y^{\prime}_{s}<0$, for all $s\geq 0$.
We are now going to show that $y^{\prime\prime}_{s}>0$, for all $s\geq 0$. Since we proved that $y_{s}^{\prime}<0$, for all $s\geq 0$, we deduce from (4.35) that
$y_{s}^{\prime\prime\prime}<\rho y_{s}^{\prime\prime},\quad\forall s\geq 0\,.$ (4.36)
Assume by contradiction that $y^{\prime\prime}_{s_{2}}\leq 0$ for some $s_{2}\geq 0$. Then, from (4.36) we get that $y^{\prime\prime\prime}_{s_{2}}<0$. It follows that there exist $\vartheta<0$, $\tau>0$ such that $y^{\prime\prime}_{s_{2}+\tau}=\vartheta<0$. Using again (4.36), we get
$\begin{dcases}y^{\prime\prime\prime}_{s}<\rho y^{\prime\prime}_{s},&\forall s\geq s_{2}+\tau,\\ y^{\prime\prime}_{s_{2}+\tau}=\vartheta.\end{dcases}$
Therefore, $y^{\prime\prime}_{s}\leq\vartheta{\mathrm{e}}^{\rho s}<0$, for all $s\geq s_{2}+\tau$. Considering that $y_{s}^{\prime}<0$, for all $s\geq 0$, this implies that the graph of $s\mapsto y_{s}$ lies below a straight line with strictly negative slope, contradicting the fact that $y_{s}>y_{\infty}$, for all $s\geq 0$, as previously established.
So, we have proved that, for all $s\geq 0$,
$y_{s}>y_{\infty},\quad y^{\prime}_{s}<0,\quad y^{\prime\prime}_{s}>0.$
This implies that there exists $\bar{y}\coloneqq\displaystyle\lim_{s\to\infty}y_{s}\geq y_{\infty}$ and that $\lim_{s\to\infty}y^{\prime}_{s}=0.$ However, it cannot be that $\bar{y}>y_{\infty}$, as it would imply, using (4.29), that $\lim_{s\to\infty}y_{s}^{\prime\prime}>0$, contradicting the fact that $y_{s}^{\prime}\to 0$, as $s\to+\infty$. Therefore, $\bar{y}=y_{\infty}$. ∎
We are now ready to establish existence and uniqueness of an equilibrium for the mean-field game.
Proposition 4.13.
For each fixed random variable $\xi$ satisfying Assumption 2.1, there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ for the mean-field game with infinite time horizon, among all equilibria verifying condition (4.3).
More precisely, $\widehat{\mathbf{u}}=z^{(\infty,\widehat{q})}$, where $z^{(\infty,q)}$ is the function defined in (3.10), and $\widehat{q}$ is the unique solution to (4.6), with initial condition $x={\mathbb{E}}[\xi]$.
Proof.
Let $\widehat{q}$ be the unique solution to (4.6), with $x={\mathbb{E}}[\xi]$. We need to check that condition (4.3) is satisfied. We have the following three cases.
Case $x=y_{\infty}$. From Theorem 4.12, we know that $\widehat{q}_{s}=y_{\infty}$, for all $s\geq 0$. Therefore,
$z^{(\infty,\widehat{q})}_{s}=K\coloneqq\dfrac{y_{\infty}^{-\beta}}{\rho+\delta},\quad\forall s\geq 0\,,$ (4.37)
and hence $\widehat{q}$ verifies condition (4.3).
Case $x>y_{\infty}$. From Theorem 4.12, we know that $\widehat{q}$ is monotone decreasing and converges to $y_{\infty}$. This implies that, $0<y_{\infty}<\widehat{q}_{s}\leq x$, for all $s\geq 0$, and that $z^{(\infty,\widehat{q})}$ is positive and bounded above by the constant $K$ defined in (4.37). Therefore, $\widehat{q}$ satisfies condition (4.3).
Case $x<y_{\infty}$. From Theorem 4.12, we know that $\widehat{q}$ is monotone increasing and converges to $y_{\infty}$. This implies that, $0<x\leq\widehat{q}_{s}<y_{\infty}$, for all $s\geq 0$, and that $z^{(\infty,\widehat{q})}$ is positive and bounded above by $\frac{x^{-\beta}}{\rho+\delta}$. Therefore, $\widehat{q}$ satisfies condition (4.3).
Applying Theorem 4.1-(iv) we get the result. ∎
Remark 4.14.
Also in this case the results of this section can be easily extended to the case where we consider an initial time time $t>0$.
Clearly, we need to adapt (in an obvious way) the definition of equilibrium given in Definition 2.3 to include the initial time $t$. The statement of Theorem 4.1 is also adapted accordingly.
The integro-differential equation (4.6) becomes
$\begin{dcases}\dfrac{{\mathrm{d}}}{{\mathrm{d}}s}y_{s}=-\delta y_{s}+\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta)(u-s)}y_{u}^{-\beta}\,{\mathrm{d}}u,&s\in[t,T],\\ y_{t}=x\,,\end{dcases}$ (4.38)
while equation (4.7) becomes
$\begin{dcases}\dfrac{{\mathrm{d}}}{{\mathrm{d}}s}z_{s}={\mathrm{e}}^{(\rho+2\delta)s}\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta-\delta\beta)u}z_{u}^{-\beta}\,{\mathrm{d}}u,&s\in[t,T],\\ z_{t}={\mathrm{e}}^{\delta t}x\,.\end{dcases}$ (4.39)
The sets in which we look for solutions to (4.39) (now dependent on $t$) are
${\mathcal{C}}_{t,x}\coloneqq\{f\in{\mathrm{C}}([t,T];(0,+\infty))\colon{\mathrm{e}}^{\delta t}\underline{\alpha}^{x}_{s}\leq f_{s}\leq{\mathrm{e}}^{\delta t}\overline{\alpha}^{x}_{s},\,\forall s\in[t,T]\},$ (4.40)
in the finite time horizon case, and
${\mathcal{C}}_{t,x,\infty}\coloneqq\{f\in{\mathrm{C}}([t,+\infty);(0,+\infty))\colon{\mathrm{e}}^{\delta t}\underline{\alpha}^{x,\infty}_{s}\leq f_{s}\leq{\mathrm{e}}^{\delta t}\overline{\alpha}^{x,\infty}_{s},\,\forall s\geq t\},$ (4.41)
in the infinite time horizon case.
The statements of Theorems 4.9 and 4.12, and of Propositions 4.10 and 4.13, remain the same (except for minor modifications). The bounds given in (4.14) become
${\mathrm{e}}^{-\delta(s-t)}x\leq y_{s}\leq{\mathrm{e}}^{-\delta(s-t)}\overline{\alpha}^{x}_{s},\qquad s\in[t,T]\,,$ (4.42)
while those given in (4.23) become
${\mathrm{e}}^{-\delta(s-t)}x\leq y_{s}\leq{\mathrm{e}}^{-\delta(s-t)}\left[x-\dfrac{x^{-\beta}}{\delta(1+\beta)(\rho+\delta-\delta\beta)}\right]+\dfrac{{\mathrm{e}}^{\delta\beta(s-t)}x^{-\beta}}{\delta(1+\beta)(\rho+\delta-\delta\beta)},\quad s\geq t.$ (4.43)
5. The deterministic model
In this section we discuss the deterministic version of the mean-field game problem introduced in Section 2. The structure of the problem will allow us to exploit the results given in Sections 3 and 4, to get the unique equilibrium of the mean-field game in an explicit form.
Let us consider the ordinary differential equation (ODE)
[…]
where $\delta>0$ is a given coefficient and the control ${\mathbf{u}}\coloneqq(u_{s})_{s\in[0,T]}$ is chosen in either of the following two classes of admissible controls: if $T<+\infty$,
${\mathcal{U}}_{T}\coloneqq$$\biggl{\{}{\mathbf{u}}\colon[0,T]\to[0,+\infty)\text{ measurable and s.t. }\int_{0}^{T}u_{s}^{2}\,{\mathrm{d}}s<+\infty\biggr{\}}\,;$ (5.2)
if, instead, $T=+\infty$,
${\mathcal{U}}_{\infty}\coloneqq$$\biggl{\{}{\mathbf{u}}\colon[0,\infty)\to[0,+\infty)\text{ measurable and s.t. }\int_{0}^{s}u_{r}\,{\mathrm{d}}r<+\infty,\,\forall s\geq 0,$
[…]
Note that, for any $x\in\mathbb{R}$ and any ${\mathbf{u}}\in{\mathcal{U}}_{T}$, Equation (5.1) has a unique solution $X^{x,{\mathbf{u}}}$, given by
$X_{s}^{x,{\mathbf{u}}}={\mathrm{e}}^{-\delta s}x+\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}u_{r}\,{\mathrm{d}}r,\quad s\in[0,T]\,.$ (5.4)
As in the stochastic case, $X$ describes the evolution of the production capacity of a representative agent, which depreciates at a rate $\delta$ and can be increased by choosing the investment rate ${\mathbf{u}}\in{\mathcal{U}}_{T}$.
Next, we consider the discounted net profit functional
$J_{T,q}(x,{\mathbf{u}})\coloneqq\int_{0}^{T}{\mathrm{e}}^{-\rho s}\left\{X_{s}^{x,{\mathbf{u}}}q_{s}^{-\beta}-\dfrac{1}{2}u_{s}^{2}\right\}{\mathrm{d}}s,$ (5.5)
where $\rho>0$ is a discount factor, $\beta>0$ is a fixed parameter, and $q=(q_{s})_{s\in[0,T]}$ is a given deterministic measurable function.
We introduce the following assumption.
Assumption 5.1.
The initial distribution of the agents in the economy $\nu_{0}$ has a density $m_{0}$ with respect to the Lebesgue measure on $\mathbb{R}$, is supported on $(0,+\infty)$, and has finite first moment. More precisely, there exists $m_{0}\in{\mathrm{L}}^{1}((0,+\infty);(0,+\infty))$ such that
$\nu_{0}(A)=\int_{A}m_{0}(x)\,{\mathrm{d}}x,\quad\forall A\in\mathscr{L}(\mathbb{R}),\quad\int_{0}^{+\infty}m_{0}(x)\,{\mathrm{d}}x=1,\quad\int_{0}^{+\infty}x\,m_{0}(x)\,{\mathrm{d}}x<+\infty\,.$
Since the representative agent is chosen randomly by picking her/his initial production capacity level $x$ according to the initial distribution $\nu_{0}$, the assumption above ensures that, for any ${\mathbf{u}}\in{\mathcal{U}}_{T}$, $X_{s}^{x,{\mathbf{u}}}\geq 0$, for all $s\in[0,T]$. Thus, the production capacity level of the representative agent is never negative.
To introduce the definition of equilibrium for the deterministic mean-field game, we need to consider the so-called continuity equation (also known as Liouville equation, cf. [5]) associated to ODE (5.1)
$\begin{dcases}\dfrac{\partial}{\partial s}p(s,x)+\dfrac{\partial}{\partial x}\left(p(s,x)[-\delta x+u_{s}]\right)=0,&(s,x)\in(0,T]\times(0,+\infty)\,,\\ p(0,x)=p_{0}(x),&x\in(0,+\infty)\,,\end{dcases}$ (5.6)
where ${\mathbf{u}}=(u_{s})_{s\in[0,T]}\in{\mathcal{U}}_{T}$ and $p_{0}$ is a given initial probability density function. This equation describes the evolution of $p_{0}$ under the flow determined by ODE (5.1).
Solutions to (5.6) are understood in the weak sense, according to the following definition.
Definition 5.2.
A function $p\in{\mathrm{L}}^{1}([0,T]\times(0,+\infty);(0,+\infty))$ is a weak solution to (5.6) if, for any function $\varphi\in{\mathrm{C}}^{\infty}([0,T]\times(0,+\infty))$ with compact support, and any $t>0$, we have
$\int_{0}^{+\infty}\varphi(t,x)p(t,x)\,{\mathrm{d}}x$ (5.7)
$\quad=\int_{0}^{+\infty}\varphi(0,x)p_{0}(x)\,{\mathrm{d}}x+\int_{0}^{t}\int_{0}^{+\infty}\left\{\dfrac{\partial}{\partial s}\varphi(s,x)+[-\delta x+u_{s}]\dfrac{\partial}{\partial x}\varphi(s,x)\right\}p(s,x)\,{\mathrm{d}}x\,{\mathrm{d}}s\,.$
Lemma 5.3.
For any ${\mathbf{u}}\in{\mathcal{U}}_{T}$ and any initial probability density $p_{0}$, supported on $(0,+\infty)$ and with finite first moment, the continuity equation (5.6) has a unique solution, given by
$p^{p_{0},{\mathbf{u}}}(s,x)={\mathrm{e}}^{\delta s}p_{0}\left({\mathrm{e}}^{\delta s}x-\int_{0}^{s}{\mathrm{e}}^{\delta r}u_{r}\,{\mathrm{d}}r\right),\quad(s,x)\in[0,T]\times(0,+\infty)\,.$ (5.8)
Moreover, for each $s\in[0,T]$, the probability density $p^{p_{0},{\mathbf{u}}}(s,\cdot)$ has finite first moment, given by
$\int_{0}^{+\infty}x\,p^{p_{0},{\mathbf{u}}}(s,x)\,{\mathrm{d}}x={\mathrm{e}}^{-\delta s}\int_{0}^{+\infty}x\,p_{0}(x)\,{\mathrm{d}}x+\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}u_{r}\,{\mathrm{d}}r\,\quad s\in[0,T]\,.$ (5.9)
Proof.
Fix ${\mathbf{u}}\in{\mathcal{U}}_{T}$ and $p_{0}$ as above. The fact that $p^{p_{0},{\mathbf{u}}}$ is a solution to (5.6) can be verified via standard computations (see, e.g., [5, Example 2.5.2] or [7, Lemma 4.15]). Uniqueness follows from an argument similar to the one used in the proof of [7, Lemma 4.16].
Finally, using the change of variables $y={\mathrm{e}}^{\delta s}x-\int_{0}^{s}{\mathrm{e}}^{\delta r}u_{r}\,{\mathrm{d}}r$, we get that, for each $s\in[0,T]$,
$\int_{0}^{+\infty}x\,p^{p_{0},{\mathbf{u}}}(s,x)\,{\mathrm{d}}x=\int_{0}^{+\infty}x\,{\mathrm{e}}^{\delta s}p_{0}\left({\mathrm{e}}^{\delta s}x-\int_{0}^{s}{\mathrm{e}}^{\delta r}u_{r}\,{\mathrm{d}}r\right){\mathrm{d}}x$
$=$ $\int_{0}^{+\infty}{\mathrm{e}}^{-\delta s}\left(y+\int_{0}^{s}{\mathrm{e}}^{\delta r}u_{r}\,{\mathrm{d}}r\right)p_{0}(y)\,{\mathrm{d}}y={\mathrm{e}}^{-\delta s}\int_{0}^{+\infty}y\,p_{0}(x)\,{\mathrm{d}}y+\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}u_{r}\,{\mathrm{d}}r\,,$
which is finite thanks to the assumptions on $p_{0}$ and ${\mathbf{u}}$. Thus, $p^{p_{0},{\mathbf{u}}}(s,\cdot)$ has finite first moment, for any $s\in[0,T]$, which satisfies (5.9). ∎
We are now ready to state the definition of equilibrium for our deterministic mean-field game.
Definition 5.4.
Fix an initial distribution $\nu_{0}$ under Assumption 5.1 and let $m_{0}$ be its density with respect to the Lebesgue measure.
A pair $(\widehat{\mathbf{u}},\widehat{q})$, where $\widehat{\mathbf{u}}\in{\mathcal{U}}_{T}$ and $\widehat{q}\colon[0,T]\to(0,+\infty)$ is a measurable function, is an equilibrium of the deterministic mean-field game if
  1. (i)
    $J_{T,\widehat{q}}(x,\widehat{\mathbf{u}})\geq J_{T,\widehat{q}}(x,{\mathbf{u}})$, for all ${\mathbf{u}}\in{\mathcal{U}}_{T}$ and $\nu_{0}$-a.e. $x>0$;
  2. (ii)
    $\widehat{q}_{s}=\int_{0}^{+\infty}x\,m^{m_{0},\widehat{\mathbf{u}}}(s,x)\,{\mathrm{d}}x$, for all $s\in[0,T]$, where $m^{m_{0},\widehat{\mathbf{u}}}$ is the unique solution of the continuity equation (5.6), given by (5.8), with initial condition $m_{0}$ and control $\widehat{\mathbf{u}}$.
Also in this case, we can adopt a fixed point argument to show that there exists a unique equilibrium, by solving, first, the problem of maximizing (5.5) for a given measurable function $q$, then determining the optimally controlled dynamics of the production capacity level of the representative agent and, finally, setting the fixed point argument from condition (ii) in Definition 5.4.
5.1. The optimization problem
Let us consider the optimization problem
$V_{T,q}(x)\coloneqq\sup_{{\mathbf{u}}\in{\mathcal{U}}_{T}}J_{T,q}(x,{\mathbf{u}}),\quad x>0\,,$ (5.10)
where $J_{T,q}$ is the discounted net profit functional defined in (5.5), for any given and fixed $q$ measurable and deterministic function.
In the finite time horizon case, i.e. $T<+\infty$, the following result holds, which is analogous to Theorem 3.3.
Theorem 5.5.
Fix $x>0$ and $q\colon[0,T]\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$. Then, $\widehat{\mathbf{u}}^{(T,q)}\coloneqq z^{(T,q)}\in{\mathcal{U}}_{T}$, where $z^{(T,q)}$ is the function defined in (3.4), is an optimal control for problem (5.10).
Moreover, $\widehat{\mathbf{u}}^{(T,q)}$ is essentially unique, i.e., if $\overline{\mathbf{u}}^{(T,q)}\in{\mathcal{U}}_{T}$ is an optimal control for problem (5.10) different from $\widehat{\mathbf{u}}^{(T,q)}$, then
$\overline{u}^{(T,q)}_{s}=\widehat{u}^{(T,q)}_{s},\quad\text{for $\mathrm{Leb}$-a.e. }s\in[0,T].$
Finally, the value function of the optimization problem admits the explicit expression
$V_{T,q}(x)=xz^{(T,q)}_{0}+\dfrac{1}{2}\int_{0}^{T}{\mathrm{e}}^{-\rho s}(z^{(T,q)}_{s})^{2}\,{\mathrm{d}}s,$ (5.11)
and the optimally controlled state $X^{x,\widehat{\mathbf{u}}^{(T,q)}}$ is given by
$X^{x,\widehat{\mathbf{u}}^{(T,q)}}_{s}={\mathrm{e}}^{-\delta s}x+\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}z^{(T,q)}_{r}\,{\mathrm{d}}r\,,\quad s\in[0,T]\,.$ (5.12)
Proof.
Replicating the argument of the proof of Proposition 3.1, we get that
$J_{T,q}(x,{\mathbf{u}})=xz^{(T,q)}_{0}+\int_{0}^{T}{\mathrm{e}}^{-\rho s}\left\{z^{(T,q)}_{s}u_{s}-\dfrac{1}{2}u_{s}^{2}\right\}{\mathrm{d}}s\,.$
Then, the result follows using the same reasoning as in the proof of Theorem 3.3. ∎
In the infinite time horizon case, i.e. $T=+\infty$, we have the following statement, which is the deterministic counterpart of Theorem 3.6. We state it without proof.
Theorem 5.6.
Fix $x>0$ and consider a function $q\colon[0,+\infty)\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}_{\rho+\delta}((0,+\infty);(0,+\infty))$ and such that the function $z^{(\infty,q)}$, defined in (3.10) is bounded on $[0,+\infty)$.
Then, $\widehat{\mathbf{u}}^{(\infty,q)}\coloneqq z^{(\infty,q)}\in{\mathcal{U}}_{\infty}$ is an optimal control for problem (5.10). Moreover, $\widehat{\mathbf{u}}^{(\infty,q)}$ is essentially unique, i.e., if $\overline{\mathbf{u}}^{(\infty,q)}\in{\mathcal{U}}_{\infty}$ is an optimal control for problem (5.10) different from $\widehat{\mathbf{u}}^{(\infty,q)}$, then
$\overline{u}^{(\infty,q)}_{s}=\widehat{u}^{(\infty,q)}_{s},\quad\text{for $\mathrm{Leb}$-a.e. }s\geq 0.$
Finally, the value function of the optimization problem admits the explicit expression
$V_{\infty,q}(x)=xz^{(\infty,q)}_{0}+\dfrac{1}{2}\int_{0}^{+\infty}{\mathrm{e}}^{-\rho s}(z^{(\infty,q)}_{s})^{2}\,{\mathrm{d}}s,$ (5.13)
and the optimally controlled state $X^{x,\widehat{\mathbf{u}}^{(\infty,q)}}$ is given by
$X^{x,\widehat{\mathbf{u}}^{(\infty,q)}}_{s}={\mathrm{e}}^{-\delta s}x+\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}z^{(\infty,q)}_{r}\,{\mathrm{d}}r\,,\quad s\geq 0\,.$ (5.14)
5.2. Existence and uniqueness of equilibria
Also in the deterministic mean-field game previously introduced, the search for an equilibrium boils down to finding a fixed point of a suitable map. According to Definition 5.4, this map is $(q_{s})_{s\in[0,T]}\mapsto\left(\int_{0}^{+\infty}x\,m^{m_{0},\widehat{\mathbf{u}}^{(T,q)}}(s,x)\,{\mathrm{d}}x\right)_{s\in[0,T]}$, where $\widehat{\mathbf{u}}^{(T,q)}$ is the optimal control for the optimization problem (5.10) – whose expression is given either in Theorem 5.5, in the case $T<+\infty$, or in Theorem 5.6, in the case $T=+\infty$ – and $m^{m_{0},\widehat{\mathbf{u}}^{(T,q)}}$ is the unique solution to (5.6), with initial probability density $m_{0}$ and control $\widehat{\mathbf{u}}^{(T,q)}$.
The next result, which is analogous to Theorem 4.1, shows that also in the deterministic case the fixed point map is the solution map of the same integral equation studied in Section 4. We omit its proof, which is a straightforward adaptation of the proof of Theorem 4.1.
Theorem 5.7.
Let us fix an initial distribution $\nu_{0}$ satisfying Assumption 5.1, and let $m_{0}$ be its density with respect to the Lebesgue measure.
Consider the deterministic mean-field game in the finite time horizon case, i.e., $T<+\infty$. Then,
  1. (i)
    If there exists an equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ in the sense of Definition 5.4, such that
    $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty)),$ (5.15)
    then $\widehat{q}$ is a solution to the integral equation
    $y_{s}={\mathrm{e}}^{-\delta s}\int_{0}^{+\infty}y\,m_{0}(y)\,{\mathrm{d}}y+\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}\int_{r}^{T}{\mathrm{e}}^{-(\rho+\delta)(u-r)}y_{u}^{-\beta}\,{\mathrm{d}}u\,{\mathrm{d}}r,\quad s\in[0,T]\,.$ (5.16)
  2. (ii)
    Vice versa, if there exist a unique solution $\widehat{q}$ to (5.16) satisfying (5.15), then there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})=(z^{(T,\widehat{q})},\widehat{q})$ of the mean-field game among all equilibria $(\widetilde{\mathbf{u}},\widetilde{q})$ such that $\widetilde{q}$ verifies (5.15), where $z^{(T,q)}$ is the function defined in (3.4).
Consider, instead, the deterministic mean-field game in the infinite time horizon case, i.e., $T=+\infty$. Then,
  1. (iii)
    If there exists an equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ in the sense of Definition 5.4, such that
    $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}_{\rho+\delta}((0,+\infty);(0,+\infty))\quad\text{and}\quad z^{(\infty,\widehat{q})}\text{ is bounded on }[0,+\infty)\,,$ (5.17)
    where $z^{(\infty,q)}$ is the function defined in (3.10), then $\widehat{q}$ is a solution to the integral equation
    $y_{s}={\mathrm{e}}^{-\delta s}\int_{0}^{+\infty}y\,m_{0}(y)\,{\mathrm{d}}y+\int_{0}^{s}{\mathrm{e}}^{-\delta(s-r)}\int_{r}^{+\infty}{\mathrm{e}}^{-(\rho+\delta)(u-r)}y_{u}^{-\beta}\,{\mathrm{d}}u\,{\mathrm{d}}r,\quad s\geq 0\,.$ (5.18)
  2. (iv)
    Vice versa, if there exist a unique solution $\widehat{q}$ to (5.18) satisfying (5.17), then there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})=(z^{(\infty,\widehat{q})},\widehat{q})$ of the mean-field game among all equilibria $(\widetilde{\mathbf{u}},\widetilde{q})$ such that $\widetilde{q}$ verifies (5.17).
Finally, we get the following result, which establishes existence and uniqueness of an equilibrium also for the deterministic mean-field game. We omit its proof, since it follows the same lines of the proofs of Propositions 4.10 and 4.13, except for minor adaptations.
Proposition 5.8.
For each fixed $\nu_{0}$ satisfying Assumption 5.1, there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ for the deterministic mean-field game with finite time horizon (resp., infinite time horizon), among all equilibria verifying the integrability condition (5.15) (resp., (5.17)).
More precisely, $\widehat{\mathbf{u}}=z^{(T,\widehat{q})}$, where $z^{(T,\widehat{q})}$ is the function defined in (3.4) (resp. (3.10)), and $\widehat{q}$ is the unique solution to (4.6), with initial condition $x=\int_{0}^{+\infty}y\,m_{0}(y)\,{\mathrm{d}}y$, where $m_{0}$ is the density of $\nu_{0}$ with respect to the Lebesgue measure.